### v1

# A quantum information theoretic analysis of reinforcement learning-assisted quantum architecture search

Abstract
In the field of quantum computing, variational quantum algorithms (VQAs) represent a pivotal category of quantum solutions across a broad spectrum of applications. These algorithms demonstrate significant potential for realising quantum computational advantage. A fundamental aspect of VQAs involves formulating expressive and efficient quantum circuits (namely ansatz) and automating the search of such ansatz is known as quantum architecture search (QAS). RL-QAS involves optimising QAS using reinforcement learning techniques. This study investigates RL-QAS for crafting ansatzes tailored to the variational quantum state diagonalization problem. Our investigation includes a comprehensive analysis of various dimensions, such as the entanglement thresholds of the resultant states, the impact of initial conditions on the performance of RL-agent, the phase change behavior of correlation in concurrence bounds, and the discrete contributions of qubits in deducing eigenvalues through conditional entropy metrics. We leverage these insights to devise an optimal, admissible QAS to diagonalize random quantum states. Furthermore, the methodologies presented herein offer a generalised framework for constructing reward functions within RL-QAS applicable to variational quantum algorithms.
I Introduction
Quantum computing represents a paradigm-shifting approach to computation, leveraging the principles of quantum mechanics to process information in ways fundamentally different from classical computers. At its core, this technology employs quantum bits, or qubits, which can exist in superposition states, enabling parallel computation paths. The information in qubits is transformed via quantum circuits consisting of quantum logic gates that are capable of acting on these computation paths in tandem. The quantum circuits represent a corresponding quantum algorithm that can effectuate a computational advantage for a target application Harrow and Montanaro (2017); Montanaro (2016); Sadhu et al. (2023). These algorithms are either constructed manually, based on conventional logical reasoning or automatically by composing the quantum gates to replicate a required input-output behaviour via training. The latter approach is termed quantum architecture search (QAS) Zhang et al. (2022); Du et al. (2022); Kuo et al. (2021); Zhang et al. (2021).
QAS typically consists of two parts. First, a template of the circuits is built, called the ansatz. The ansatz can have parametric quantum gates, e.g., rotation angles. Then, these parameters are determined via the variational principle using a classical optimiser in a feedback loop. Algorithms constructed via this technique are called variational quantum algorithms (VQA) McClean et al. (2016); Cerezo et al. (2021). QAS can also used to determine non-parametric circuits as an approach for quantum program synthesis Sarkar (2022). In this research, however, we will focus on QAS in the context of VQA.
The ansatz design is critical, as it directly influences the expressivity and efficiency of the quantum solution. Given the vast potential configuration space of quantum circuits, the process of identifying a suitable ansatz is a challenging problem. In previous works such as Ostaszewski et al. (2021); Kundu et al. (2024); Patel et al. (2024); Du et al. (2022), the search for the quantum circuits that solve various VQA has been automated using reinforcement learning (RL) techniques Sutton and Barto (2018). We particularly call quantum ansatzes capable of expressing the solution of VQA an admissible ansatz. The automated search for such admissible ansatz using the RL-agent is termed reinforcement learning-assisted quantum architecture search (RL-QAS). In this setting, after a certain number of episodes, the RL agent returns possible configurations of quantum ansatz that solve the problem. For $E_{tot}$ episodes, the agent returns $E_{s}$ successful episodes, and in $E_{tot}-E_{s}$ episodes, the agent fails to find a configuration within a specific predefined accuracy and depth. It has been shown in Ostaszewski et al. (2021); Wu et al. (2023); Kundu et al. (2024); Patel et al. (2024); Kundu (2024) that classical machine learning-driven QAS algorithms outperform the state-of-art structures of ansatz and effectively optimise the number of parameters, depth, and impact of noise in VQAs.
The use-cases of such admissible ansatzes cover applications in quantum information theory LaRose et al. (2019); Cerezo et al. (2021); Tan and Volkoff (2021); Cerezo et al. (2020); Kundu and Miszczak (2022), chemistry Kandala et al. (2017); Mustafa et al. (2022); Delgado et al. (2021) and combinatorial optimization Khairy et al. (2020); Liu et al. (2022); Glos et al. (2022). Yet, a thorough investigation of such admissible ansatzes is still missing in recent literature. It has been previously shown Patel et al. (2024) that in RL-QAS methods, it is crucial to feed proper information about the problem to the agent to enhance the agent’s performance. One of the ways of providing this information is by engineering a reward function. In RL, the formulation of the reward function and its engineering can be sparse and dense. The choice of optimal reward function depends on trial and error by investigating the performance of the agent under various reward configurations.
In this article, we investigate the admissible ansatzes proposed by the RL-agent from the perspective of quantum information theory Wilde (2013). For the investigation, we specifically focus on the reinforcement learning enhanced variational quantum state diagonalisation (RL-VQSD) problem Kundu et al. (2024). In the VQSD algorithm, the problem-inspired ansatz is based on density matrix exponentiation Lloyd et al. (2014), but its depth and number of gates increase exponentially with the number of qubits. Hence, searching for an optimal ansatz (with a minimum number of gates and depth) that diagonalises a quantum state is an open research problem. In RL-VQSD, the search for the optimal configuration of ansatz is done by using a double deep-Q network (DDQN). See Appendix B for further details on the VQSD algorithm, the agent and environment specifications. The RL-agent proposes $E_{s}<E_{tot}$ admissible ansatzes that solves the VQSD problem.
Our investigation shows that the concurrence of the $s$ ansatzes lies between an upper and lower bound, independent of the different initialisation of the double deep-Q network (DDQN), where each initialisation of DDQN corresponds to a different random initialisation of the weights. Moreover, we show that the upper and lower bounds are anti-correlated with respect to the increment in the initial state entanglement. The anti-correlation turns into a mild, and later strong, correlation when the initial state entanglement goes beyond $0.322$, indicating a phase change in correlation between the upper and lower bounds of the concurrence. The phase transition provides us with an in-depth insight into the relation between the input and the output of the VQSD algorithm. Further, our investigation reveals that the RL-agent can generate high concurrence admissible ansatz with fewer 2-qubit gates and circuit depth compared to ansatzes with very low concurrence. Furthermore, we quantify the contribution of individual qubits of the RL-ansatz using conditional quantum entropy. Using such a measure, we observe a correlation between entanglement and quantum entropy, revealing why RL-VQSD can efficiently find the largest eigenvalues but fails to find the smallest ones.
The structure of the paper is as follows. We present an overview of our main results in Sec. II. In Sec. III.1, we analyse the upper and lower bounds on the concurrence of the admissible ansatzes that solve the VQSD problem. Specifically, in Sec. III.1.1 we present the variation of the entanglement bounds for different weight initialisation of the DDQN. We observe the correlation properties of the entanglement bounds dependent on the concurrence of the initial state in Sec. III.1.2. In Sec. III.1.3, we present the dependence of the entanglement bounds on the number of gates and circuit depth of the ansatz. The analysis of the contribution of individual qubits to the admissible ansatz is provided in Sec. III.2. In particular, we present the contribution of each qubit for change in the entanglement of the ansatz in Sec. III.2.1. In Sec. III.2.2, we explain why RL-VQSD can efficiently find the largest eigenvalues of the initial state but not the smallest eigenvalues. We provide concluding remarks and discuss possible future directions in Sec. IV.
II Contributions
Through the following points, we summarise the contribution of the paper.
  1. 1.
    Investigating admissible ansatzes of reinforcement learning based quantum architecture search problem from a quantum information theory perspective. We focus on the RL-VQSD algorithm, where the proposed ansatzes diagonalise random quantum states.
  2. 2.
    Demonstrating that the entanglement of the states generated by the RL-agent-proposed admissible ansatzes lies within a lower and upper concurrence bound.
  3. 3.
    The upper and lower bounds of concurrence remain consistent across various weight initialisation of the DDQN.
  4. 4.
    Bounds are anti-correlated with respect to increasing entanglement until concurrence surpasses $0.322$; after this point, the anti-correlation turns into correlation. Hence, we introduce the concurrence $0.322$ as the phase change point.
  5. 5.
    Investigation of ansatzes in upper and lower bounds, showing optimal ansatz for VQSD generates high concurrence states.
  6. 6.
    RL-agent requires fewer gates and circuit depth to produce highly entangled admissible ansatzes than weakly entangled ansatzes. We utilise this information to minimise RL-agent and environment (which is the configuration of ansatz) interaction, thereby optimising the convergence time of the RL-VQSD.
  7. 7.
    We evaluate individual qubits’ contribution in diagonalising 2-qubit quantum states using conditional quantum entropy.
  8. 8.
    Our investigation shows that the qubits have equal contributions on average in diagonalising random quantum states.
  9. 9.
    We further investigate the correlation in conditional entropy between qubits for different inferred eigenvalues. It requires mild anti-correlation for the first two largest eigenvalues and strong anti-correlation for the smallest eigenvalues.
  10. 10.
    The majority of admissible ansatzes lie in a mild anti-correlation regime, explaining VQSD’s limitation in finding the smallest eigenvalues.
III Main results
In this section, we investigate the upper and lower bounds of concurrence of the ansatzes produced by the RL agent in diagonalising $2$-qubit random quantum states sampled from IBM qiskit’s random_density_matrix module. For the simulation, we consider multiple such quantum states (specified in the caption of figures), and for each state, we use $10000$ episodes of RL-VQSD to diagonalise the state. The configuration of the RL-VQSD is provided in detail in Appendix B.
III.1 Entanglement
III.1.1 DDQN weight independent entanglement bound
To benchmark the performance of the RL agent for a specific problem, it is ideal to take multiple initialisation of the DDQN (see Appendix B for a brief description).
state max. concurrence min. concurrence
no. avg. std. div avg. std. div.
1 0.874 0.002 0.183 0.004
2 0.849 0.002 0.108 0.028
3 0.975 0.001 0.084 0.006
4 0.936 0.002 0.118 0.009
5 0.880 0.001 0.188 0.013
6 0.911 0.003 0.156 0.024
7 0.933 0.001 0.235 0.013
8 0.928 0.005 0.208 0.013
9 0.831 0.011 0.302 0.010
Table 1: In this table, we present the average and standard deviation of the concurrence of the final state after evolution through the circuit for five different initialisations of the DDQN wrights for different quantum seed states.
To efficiently investigate the upper and lower bounds of concurrence Hill and Wootters (1997); Wootters (1998) produced by the RL ansatz, we first sample different random quantum states, then diagonalize them using RL-VQSD Kundu et al. (2024) for five different initialisation of DDQN weights. Thereafter, the agent suggests different possible configurations for the RL-ansatz, which diagonalizes the sampled random quantum states. For a specific state, we collect such valid quantum circuits and find the concurrence of the state obtained after evolving through the RL-ansatz. It should be noted that finding the concurrence of the evolved state indicates how much the RL-ansatz induces entanglement to the input sampled state. See Appendix C for further analysis of the sampled random quantum states.
The results are presented in Table 1. We observe that the optimal configuration of the RL-ansatz always gives us a state with concurrence close to one. Hence, the deviation in the maximum concurrence over different weights of the DDQN is more significant than the deviation in the minima of the concurrence. As seen in Table 1, the deviation in the max concurrence is in the order of $10^{-3}$. The low standard deviation implies that there is a negligible deviation in concurrence across different initialisation of the DDQN for a specific state. This indicates that the bounds on concurrence are invariant for different configurations of the RL-ansatz. Hence, proving that the entanglement bounds of the RL-ansatz are independent of initialisation of DDQNs. Due to this observation, it is sufficient to consider a single initialisation of DDQN to benchmark the outcomes in the remaining simulations.
In the following subsection, we investigate how the upper and lower bounds on the entanglement of the RL-ansatz depend on the entanglement of the input quantum state.
III.1.2 Anti-correlation between entanglement bounds
Figure 1: Reciprocal behaviour in the upper and lower bounds of concurrence with respect to the increasing concurrence in the input random state. A clear anti-correlation between entanglement minima and maxima which breaks down after $0.3$. From the interpolating function, the Pearson correlation coefficient (PCC) for the first eleven points is -0.982, while that for the remaining points is 0.049. At this point, the correlation after the first eleven points changes from -0.055 to 0.049. There is a phase transition from weak anti-correlation to weak correlation.
Corresponding to a predefined threshold, the agent proposes different architectures of quantum circuits that diagonalise the quantum state. To observe the amount of entanglement generated by the RL-ansatz we calculate the concurrence of the state after passing through the RL-ansatz, just before the dephasing operation. By sorting out the quantum circuit with maximum and minimum entanglement, we observe their variation with the entanglement of the input state. It has been proved in the previous section that the upper and lower bounds of entanglement are independent of different initialisation of DDQNs; hence, we for further benhamarking we can constrain ourselves to just one initialisation of the DDQN weights.
In Figure 1, we illustrate the variation of the entanglement bounds with the entanglement of the input quantum state. To investigate the dependency, we use the Pearson correlation coefficient (PCC) Lee Rodgers and Nicewander (1988) as a quantifier of correlation. See Appendix A for a brief description. We see that the upper and lower bounds of entanglement are initially strongly anti-correlation when the input state entanglement is below $0.3$ and then strongly correlate afterwards, indicating a phase change in the upper and lower bounds of entanglement.
Figure 2: The PCC indicating Pearson correlation coefficient is calculated over an interval $i<k<j$. The coefficient is described as in Equation 1, where $i$ is minimum, $j$ is maximum concurrence, and $k$ splits the intermediate values of concurrence into two parts. Hence, $i$ and $j$ are fixed. The $w_{ik}\in[0,1]$ is the weight of the first interval which defines how much each of the two intervals contributes when varying $k$. We vary the cardinality of the two intervals, causing variation in the weights. In the figure, we consider $i=0.119$ and $j=0.402$ and when $k=0.322$, i.e. $i<x\leq 0.322$, the weight corresponds to the first interval.
Next, we obtain the variation of PCC between the upper and lower bound on the concurrence of the initial state. Specifically, we introduce a parameter $k\in(i,j)$ to split the range of initial state concurrence $\mathbbm{u}\coloneqq[i,j]$ into two intervals: $\mathbbm{u}_{1}\coloneqq[i,k)$ and $\mathbbm{u}_{2}\coloneqq[k,j]$. Also, let us denote the set of lower and upper bounds on the concurrence of the ansatz as $u$ and $l$, respectively. For both the ranges $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$, we introduce the following correlation function:
$\eta^{k}_{w_{ik}}=w_{ik}\textrm{PCC}^{ik}_{ul}+(1-w_{ik})\textrm{PCC}^{kj}_{ul},$ (1)
where $w_{ik}\in\{0,1\}$ and $\textrm{PCC}^{ik}_{ul}$, $\textrm{PCC}^{kj}_{ul}$ are the value of PCC between the upper and lower bounds of concurrence of the ansatz for the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$ respectively.
Note 1.
[…]
We present in Figure 2, the variation of $\eta_{0}^{k}$ (in blue) and $\eta_{1}^{k}$ (in red) for different values of $k$. We observe that $\eta_{1}^{k}\approx-1~{}\forall k\in(i,j)$. This indicates strong cumulative anti-correlation between the upper and lower bounds of concurrence. For $\eta_{0}^{k}$, we observe a gradual change of $\text{PCC}^{kj}_{ul}$ from -0.582 to 1. We note that $\eta_{0}^{k}$ changes from $-0.0451$ to $0.279$ for $k\in[0.313,0.379]$ with $\eta_{0}^{k}=0$ at $k=k_{\ast}=0.322$. We call this point the phase transition point of correlation between upper and lower bounds of concurrence for the VQSD ansatz. At values of $k<k_{\ast}$, there is anti-correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$. While for values of $k>k_{\ast}$, there is (anti-)correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $(\mathbbm{u}_{2})~{}\mathbbm{u}_{1}$.
III.1.3 Agent and environment interaction reduction based on entanglement bound
Here, we provide a detailed investigation of the admissible ansatzes produced by the RL-agent in terms of ansatz depth, the total number of 2-qubit gates and the total number of gates in the upper and lower bounds of concurrence.
Figure 3: We observe that the RL-agent can generate ansatz with very high concurrence with less number of 2-qubit gates and circuit depth compared to the ansatzes with very low concurrence. As the main aim of RL-VQSD is to diagonalise unitary with a small number of gates and depth, our analysis shows that if we particularly focus on admissible ansatzes near the upper bound of concurrence, we can not only reduce the computational cost of RL-VQSD severely but find optimal admissible ansatzes. The numbers in the bars correspond to the amount of entanglement generated by the RL-ansatz when initialised in the vacuum state.
Our investigation reveals that the admissible ansatzes corresponding to the upper bound of concurrence has, on average, a shorter depth of $13.23$ and a smaller number of gates of $18.5$ compared to the admissible ansatzes corresponding to the lower bound of concurrence where the depth is $13.9$ and number of gates is $19$ over $25$ random quantum states. Moreover, Figure 3 makes it more prominent that the minimum number of gates and the depth of the admissible ansatzes are smaller for the upper bound of concurrence than the lower. This observation is crucial in reducing the subspace of all possible admissible ansatzes. The main motivation of RL-QAS in VQA is to find ansatzes that solve a VQA with a minimum number of gates and circuit depth; in the case of RL-VQSD, we can narrow down the search for an optimal admissible ansatz near the upper bound of entanglement by encouraging the RL-anent to produce ansatzes with very high concurrence and penalise for the other structures of ansatz.
III.2 Conditional quantum entropy
III.2.1 Contribution of individual qubits for change in entanglement
We quantify the contribution of each qubit using the conditional quantum entropy $S_{q_{0}|q_{1}}^{\rho^{\prime}_{q_{0}q_{1}}}$ Cerf and Adami (1997, 1999); Wilde (2013). The conditional entropy of an individual qubit for an $N$ qubit quantum system indicates the amount of uncertainty remaining about that qubit after measuring the whole system. Hence, it is essential to know how much information the state of qubits provides about the state of the $N$ qubits. For a detailed discussion and notation, see Appendix A.
Figure 4: Relative change in the contribution of qubits 0 and 1 as a function of change in entanglement. Weight of the positive part $5.524$ and that of the negative part is $-4.854$.
Let the change in the conditional entropy of the qubit $q_{0}(q_{1})$ after evolving through the VQSD ansatz be given by $\Delta S^{\rho\to{\rho^{\prime}}}_{q_{0}|q_{1}}(\Delta S^{\rho\to{\rho^{\prime}}}_{q_{1}|q_{0}})$. We quantify the relative contribution of individual qubits via the quantity:
$\Delta_{q_{0}}^{q_{1}}\coloneqq\Delta S^{\rho\to{\rho^{\prime}}}_{q_{0}|q_{1}}-\Delta S^{\rho\to{\rho^{\prime}}}_{q_{1}|q_{0}}.$ (2)
We note that $\Delta_{q_{0}}^{q_{1}}>0$ implies a greater change in conditional entropy of $q_{0}$ as compared to $q_{1}$, and hence the ansatz has a greater contribution from $q_{0}$ as compared to $q_{1}$. Following a similar reasoning $\Delta_{q_{0}}^{q_{1}}<0$ implies $q_{1}$ has greater contribution as compared to $q_{0}$, while $\Delta_{q_{0}}^{q_{1}}=0$ implies both $q_{0}$ and $q_{1}$ contributes equally to the ansatz.
We present the variation of $\Delta_{q_{0}}^{q_{1}}$ for the change in concurrence due to the ansatz given by $\Delta_{c}\coloneqq C^{\rho_{q_{0}q_{1}}}-C^{\rho^{\prime}_{q_{0}q_{1}}}$ in Figure 4. We observe that the distribution of $\Delta_{q_{0}}^{q_{1}}$ is bounded in the interval $[-0.7,0.7]$ and is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$. This implies an equal contribution to the VQSD ansatz from both the qubits for the range of change in entanglement. In the region of $\Delta_{q_{0}}^{q_{1}}>0$, the relative contribution of $q_{0}$ first increases and then decreases. However, no such pattern is observed for the region of $\Delta_{q_{0}}^{q_{1}}<0$. When $\Delta_{c}>0.4$, the value of $\Delta_{q_{0}}^{q_{1}}$ monotonically decreases with increase in $\Delta_{c}$. Also, for $\Delta_{c}<0.2$, $\Delta_{q_{0}}^{q_{1}}$ increases monotonically with increase in $\Delta_{c}$.
III.2.2 Problem with variational quantum state diagonalization
In this section, we investigate the contribution of individual qubits to the performance of the RL-VQSD algorithm. Through this investigation, we get a deeper insight into why the variational quantum state diagonalization algorithm can not find the smallest eigenvalues.
We present the main observations in Figure 5, where we plot the four eigenvalues of $21$, $2$-qubit Haar random mixed quantum states with the correlation in conditional entropy among the qubits. For an individual quantum state, the RL-agent proposes $\mathbf{s}$ different structures of admissible ansatzes. Then we get $\rho_{i}^{\prime}$ states where $i\in[0,\mathbf{s}]$. Afterwards, we calculate the conditional entropy of individual qubits for each state and take the median. [^1] Calculating the median over the mean is justified by the broad range of conditional entropy distribution, which makes the median a more appropriate measure of central tendency. Then, we calculate the correction in conditional entropy among the qubits using PCC.
In Figure 5, we observe that a mild anti-correlation between the qubits is required to find the two largest eigenvalues. As the magnitude of the largest eigenvalues decreases, the mild anti-correlation turns into a strong correlation between the qubits. This indicates a strong correlation between the qubits is required to achieve the two largest eigenvalues. Meanwhile, we require a strong anti-correlation between the qubits to find the smallest eigenvalues of the same states. Then, as the magnitude of the smallest eigenvalue increases, the strong anti-correlation turns into mild anti-correlation.
Figure 5: The conditional entropy correlation between the first and the second qubit decreases as the magnitude of the largest and the 2nd largest eigenvalues increases. Meanwhile, the same correlation between the qubits increases with the magnitude of the 3rd and the smallest eigenvalues. The RL-ansatz proposes specific structures of admissible ansatzes, and for each such ansatz, it is not possible to have two different kinds of correlations. Hence, our investigation suggests that it is not feasible to find the smallest eigenvalues with the same ansatz that provides a good approximation of the largest eigenvalues. The correlation of conditional entropy between qubits is calculated using Pearson correlation function (PCC) and is the median over all the configuration of RL-ansatz proposed by the agent. The $\bigstar$ signifies the eigenvalues corresponding to the same state.
We recall that the correlation is calculated among the qubits based on the conditional entropy of individual qubits. As it is not possible to have two different kinds (such as mild anti-correlation and strong anti-correlation at the same time) of correlation between the qubits in the same ansatz structure, we can not find the largest and smallest eigenvalues utilising the same ansatz. For example, we mark the eigenvalues of a state with $\bigstar$, where we observe that to find the largest two eigenvalues, the RL-ansatz need to induce a mild anti-correlation between the two qubits; meanwhile, to obtain the smallest eigenvalues, we require strong anti-correlation.
On the contrary, for the algorithm, as the agent adds a new gate to the RL-ansatz, the correlation in conditional entropy among the qubits increases. That is why it is easier for the RL-ansatz to propose admissible ansatz with mild anti-correlation. We further certify this observation in Figure 6, where we observe the correlation in conditional entropy between qubits for $21$ random quantum states. We see that most of the admissible ansatz provide a mild anti-correlation among the qubits, indicating that the VQSD algorithm primarily focuses on finding the largest eigenvalues rather than the smallest ones. Hence, it is easier for the variational diagonalisation algorithm to find the largest eigenvalues more efficiently than the smallest ones.
Figure 6: The correlation in conditional entropy among qubits is in the range of mild correlation for $21$ random quantum states. This observation, along with the observation in Figure 5 certify that RL-agent only proposes a particular configuration of admissible ansatz that can find only the largest eigenvalues. The average correlation among all the $21$ states is $-0.4263$, which is in the regime of mild anti-correlation.
IV Discussion
In this article, we use tools from quantum information theory to analyse the admissible ansatzes proposed by the RL-agent for solving reinforcement learning-based quantum architecture search problems, and we particularly investigate the RL-agent proposed ansatzes in solving recently proposed RL-VQSD problem.
We observe that the concurrence of the admissible ansatz ranges between an upper and lower bound, which are independent of the initialisation of the weights of the deep-Q network. The upper and lower bounds are initially anti-correlated with respect to the initial state entanglement. The anti-correlation turns into a mild and eventually strong correlation as the entanglement of the initial state surpasses beyond a phase change point at concurrence $0.322$. We also see that the optimal configuration of the admissible ansatz lies in the upper bound of concurrence, which has a smaller circuit depth and requires fewer gates than the lower bound. These observations provide insight regarding the relation between the entanglement of the initial state and the entanglement generated by the RL-VQSD admissible ansatzes. We effectively utilise these observations to greatly reduce the RL-agent and environment interaction, i.e., the computational time needed to solve the RL-VQSD problem. Additionally, the admissible ansatzes not only diagonalize the random quantum states with high accuracy but can be used to generate close to maximum entangled states from the vacuum state.
Furthermore, we quantify the contribution of each qubit in the admissible ansatz using conditional quantum entropy. For the admissible RL-VQSD ansatz, we observe an equal contribution from each qubit for the full range of change in the entanglement of the state due to the ansatz. Focusing on the task of obtaining the eigenvalues of the starting state, we observe that a strong correlation between the qubits of the admissible ansatz is required to obtain the two largest eigenvalues, while strong anti-correlation between the qubits is needed to find the smallest eigenvalues of the same state. Noting that it is impossible to have both strong correlation and anti-correlation between the qubits of the same ansatz, we conclude that we cannot simultaneously obtain both the largest and smallest eigenvalues from the RL-VQSD admissible ansatz. From the observation, most of the admissible ansatz has mild anti-correlation among the qubits, which explains why it is easier for the diagonalisation algorithm to find the largest eigenvalues than the smallest ones.
For future work, it would be interesting to see the implication of introducing entanglement Horodecki et al. (2009) or nonlocality Brunner et al. (2014); Sadhu and Das (2023)-based reward functions of the RL-agent for solving the VQSD problem. Specifically, given that most of the admissible ansatzes have high values of concurrence, it is expected that an entanglement or non-locality-based reward function may further reduce the time required to find the optimal configurations. Also, we may expect this to provide an optimal quantum circuit that provides both the lowest and highest eigenvalues of the input quantum state.
Acknowledgements.
A. Sadhu acknowledges the PhD fellowship from the Raman Research Institute, Bangalore. AK would like to thank Mateusz Ostaszewski for the valuable discussions. This research was partially supported by PL-Grid Infrastructure Grant Number PLG/2023/016205.
Appendix A Preliminaries
We assume that every quantum system say $S$ has an associated Hilbert space $\mathscr{H}_{S}$ of finite dimension $d_{S}$. The state of a quantum system $S$ is represented by a density operator $\rho_{S}$ defined on $\mathscr{H}_{S}$ and satisfies the conditions (a) Positivity: $\rho_{S}\geq 0$, (b) Hermitian: $\rho_{S}=\rho_{S}^{\dagger}$ and (c) Trace Preserving: $\operatorname{Tr}[\rho_{S}]=1$. We denote the set of density operators of $S$ as $\mathscr{D}(\mathscr{H}_{S})$. A pure state is a rank-one density operator given by $\rho_{S}\coloneqq\outerproduct{\psi}{\psi}_{A}$, where |$\ket{\psi}_{S}\in\mathscr{H}_{S}$ is a unit-norm vector in $\mathscr{H}_{S}$. We denote the density operator of a composite system $S_{1}S_{2}$ as $\rho_{S_{1}S_{2}}\in\mathscr{H}_{S_{1}S_{2}}$ with $\operatorname{Tr}_{S_{2}}\rho_{S_{1}S_{2}}=\rho_{S_{1}}\in\mathscr{H}_{S_{1}}$ being the reduced state of $S_{1}$. The state $\rho_{S_{1}S_{2}}\in\mathscr{H}_{S_{1}S_{2}}$ is called separable if it cannot be expressed in the form of $\rho_{S_{1}S_{2}}=\sum_{x}p_{x}\rho_{S_{1}}^{x}\otimes\rho_{S_{2}}^{x}$, where $\{\rho_{S_{1}}^{x}\}_{x}$ and $\{\rho_{S_{2}}^{x}\}_{x}$ are sets of pure states, $p_{x}\in[0,1]$ and $\sum_{x}p_{x}=1$. The states that cannot be expressed in the above form are said to be entangled Horodecki et al. (2009). Next, we briefly discuss a relevant measure of entanglement for this work.
A.1 Concurrence
We quantify the amount of entanglement present in a bipartite quantum state $\rho_{AB}\in\mathscr{D}(\mathscr{H}_{A}\otimes\mathscr{H}_{B})$ using concurrence Hill and Wootters (1997); Wootters (1998) which is defined as:
$C^{\rho_{AB}}=\max\{0,\lambda^{e}_{1}-\lambda^{e}_{2}-\lambda^{e}_{3}-\lambda^{e}_{4}\}$ (3)
where $\lambda^{e}_{s}$ are the square root of the eigenvalues of $\rho_{AB}\tilde{\rho}_{AB}$ in descending order with $\tilde{\rho}_{AB}=(\sigma_{y}\otimes\sigma_{y})\rho^{*}_{AB}(\sigma_{y}\otimes\sigma_{y})$ being the spin-flipped state of $\rho_{AB}$ and $\sigma_{y}$ being a Pauli spin matrix. For the state $\rho_{AB}=\outerproduct{\psi}{\psi}_{AB}$, the concurrence is defined as
$\mathcal{C}^{\ket{\psi}_{AB}}=\sqrt{2[1-\operatorname{Tr}\rho_{A}^{2}]},$ (4)
where $\rho_{A}=\operatorname{Tr}_{B}\rho_{AB}$ is a subsystem of the combined state. It is easy to see that the concurrence of a pure bipartite qubit maximally entangled state is one, while that of a bipartite qubit separable state is zero. In the following subsection, we discuss a measure to quantify the contribution of individual qubits in a quantum circuit.
A.2 Conditional quantum entropy
We quantify the contribution of each qubit in a quantum circuit using the conditional quantum entropy Cerf and Adami (1997, 1999), which is defined as follows.
Definition 1 (Conditional quantum entropy Wilde (2013)).
For the state $\rho^{\prime}_{AB}\in\mathscr{D}(\mathscr{H}_{A}\otimes\mathscr{H}_{B})$, the conditional quantum entropy $S_{A|B}^{\rho^{\prime}_{AB}}$ is equal to the difference of the joint quantum entropy $S^{\rho^{\prime}_{AB}}_{AB}$ and the marginal entropy $S^{\rho^{\prime}_{AB}}_{B}$:
$S_{A|B}^{\rho^{\prime}_{AB}}\coloneqq S^{\rho^{\prime}_{AB}}_{AB}-S^{\rho^{\prime}_{AB}}_{B}.$ (5)
Operationally, the conditional entropy $S_{A|B}^{\rho^{\prime}_{AB}}$ of the state for the state $\rho^{\prime}_{AB}$ signifies how mixed the marginal state $\rho^{\prime}_{A}$ ($\rho^{\prime}_{B}$) is as compared $\rho^{\prime}_{AB}$. For more details of the following theorems and observations, please refer Wilde (2013).
Observation 1.
For the state $\rho^{\prime}_{AB}\in\mathscr{D}(\mathscr{H}_{A}\otimes\mathscr{H}_{B})$, consider a purification |$\ket{\psi}_{ABE}\in\mathscr{H}_{A}\otimes\mathscr{H}_{B}\mathscr{H}_{E}$. It then follows that:
$S_{A|B}^{\rho^{\prime}_{AB}}=S_{E}^{\psi}-S_{B}^{\psi}.$ (6)
The conditional quantum information thus measures the difference in the entropy of the states $\rho_{E}\in\mathscr{D}(\mathscr{H}_{E})$ and $\rho_{B}\in\mathscr{D}(\mathscr{H}_{B})$.
Theorem 1.
For the state $\rho^{\prime}_{AB}\in\mathscr{D}(\mathscr{H}_{A}\otimes\mathscr{H}_{B})$, the following inequality applies to the marginal entropy $S_{A}^{\rho^{\prime}_{AB}}$ and the conditional quantum entropy $S_{A|B}^{\rho^{\prime}_{AB}}$:
$S_{A}^{\rho^{\prime}_{AB}}\geq S_{A|B}^{\rho^{\prime}_{AB}}.$ (7)
The above theorem states that conditioning cannot increase entropy. The value of conditional entropy can be positive or negative depending on the quantum state for which it is evaluated. However, it cannot be arbitrarily large or arbitrarily small. The following theorem places an upper bound on the absolute value of the conditional quantum entropy.
Theorem 2.
The conditional entropy is bounded as follows:
$|S_{A|B}^{\rho^{\prime}_{AB}}|\leq\log\dim(\mathscr{H}_{A})$ (8)
The bound is saturated for $\rho^{\prime}_{AB}=\mathbbm{1}_{A}\otimes\sigma_{B}$, where $\mathbbm{1}_{A}$ is the maximally mixed state and $\sigma_{B}\in\mathscr{D}(\mathscr{H}_{B})$, and also for the family of maximally entangled states.
In the following subsection, we discuss a measure to quantify the amount of correlation between two random variables.
A.3 Correlation coefficient
We quantify the amount of correlation between two random variables $r_{1}$ and $r_{2}$ using the Pearson correlation coefficient (PCCLee Rodgers and Nicewander (1988) defined as:
$\textrm{PCC}_{r_{1}r_{2}}=\frac{\mathrm{E}[(r_{1}-\mu_{r_{1}})(r_{2}-\mu_{r_{2}})]}{\sigma_{r_{1}}\sigma_{r_{2}}},$ (9)
where $\mu_{r_{j}}$, $\sigma_{r_{j}}$ are the mean and standard deviation of the distribution $r_{j}$ with $j\in\{1,2\}$ and $\mathrm{E}[.]$ is an operator that returns the expectation value of a random variable. Note that the Pearson correlation coefficient (PCC) is symmetric: $\textrm{PCC}_{r_{1}r_{2}}=\textrm{PCC}_{r_{2}r_{1}}$. The value of PCC ranges in between $-1$ to $+1$. A correlation value of +1(1)11+1(-1)+ 1 ( - 1 ) implies perfect correlation (anti-correlation) between the two random variables $r_{1}$ and $r_{2}$, while zero implies no linear dependence between the variables.
Appendix B The reinforcement learning enhanced variational quantum state diagonalisation
The VQSD algorithm LaRose et al. (2019) for a quantum state $\rho$ comprises three subroutines: (1) In the training subroutine, the parameters $\vec{\theta}$ of a quantum gate sequence $U(\vec{\theta})$ are optimised for a given state $\rho$. Ideally, after optimisation, the sequence $U(\vec{\theta}_{\text{opt}})$ satisfies
$\rho^{\prime}=U(\vec{\theta}_{\text{opt}})\rho U(\vec{\theta}_{\text{opt}})^{\dagger}=\rho_{\text{diag}},$
where $\rho_{\text{diag}}$ is $\rho$ diagonalized in its eigenbasis and $\vec{\theta}_{\text{opt}}$ are the optimal angles. Classical gradient-based methods such as SPSA and Gradient Descent, or gradient-free optimisation Powell (1994, 2006), can be utilized in the training process. (2) In the eigenvalue readout subroutine, using the optimised unitary $U(\vec{\theta}_{\text{opt}})$ and one copy of state $\rho$, one can extract all the eigenvalues for low-rank states or the largest eigenvalues for full-rank states. This is achieved by measuring $\rho^{\prime}$ in the computational basis, $\mathbf{b}=b_{1}b_{2}\ldots b_{n}$, as follows
$\lambda^{\prime}=\bra{\mathbf{b}}\rho^{\prime}\ket{\mathbf{b}},$
where $\lambda^{\prime}$ are inferred eigenvalues. (3) In the final step, we prepare the eigenvectors associated with the largest eigenvalues. If $\mathbf{b}^{\prime}$ is a bit string associated with $\lambda^{\prime}$, then the inferred eigenvectors $|v_{\mathbf{b}^{\prime}}^{\prime}\rangle$ are obtained as follows
$|{v}^{\prime}_{{\mathbf{b}}^{\prime}}\rangle=U(\theta_{\textrm{opt}})^{\dagger}\ket{{\mathbf{b}}^{\prime}}=U(\theta_{\textrm{opt}})^{\dagger}\left(X^{b_{1}}\otimes\ldots\otimes X^{b_{n}}\right)\ket{\mathbf{0}}.$
In the reinforcement learning enhanced variational quantum state diagonalisation, an RL-agent is utilised to find the optimal configuration of $U(\vec{\theta})$. The agent contains a double deep-Q network (DDQN) Van Hasselt et al. (2016) with $\epsilon$-greedy policy. The DDQN optimises the stability of Q-learning by using two networks to decouple the action selection $a^{\prime}$ in the subsequent state $s^{\prime}$ from its current reward $r$ evaluation, leading to less overestimated Q-values and more stable and reliable learning outcomes.
The key update equation is
$Q(s,a;\theta)\leftarrow Q(s,a;\theta)\\ +\alpha[r+\gamma Q(s^{\prime},\arg\max_{a^{\prime}}Q(s^{\prime},a^{\prime};\theta);\theta^{\prime})-Q(s,a;\theta)]$ (10)
where, $Q(s,a;\theta)$ is the predicted Q-value for state $s$ and action $a$, parameterised by weights $\theta$ of the online network. $\alpha$ is the learning rate and $\gamma$ is the discount factor. The settings of the parameters are similar to the ref. Kundu et al. (2024).
The optimal policy is obtained by optimising the weights of the DDQN using the ADAM optimiser. In the RL setup, the RL-state encodes the ansatz and is defined by a tensor-based encoding presented introduced in Patel et al. (2024). After each environment and agent interaction, the reward function
$R=\left\{\begin{array}[]{ll}+\mathcal{R}&\text{for }C_{t}(\vec{\theta})<\zeta+10^{-5}\\ -\textrm{log}\left(C_{t}(\vec{\theta})-\zeta\right)&\text{for }C_{t}(\vec{\theta})>\zeta\end{array}\right.,$ (11)
is calculated. The goal of the RL-agent is to reach the minimum error for a predefined threshold $\zeta$ (which is a hyperparameter of the model) by optimising the parameters of the $U(\vec{\theta})$ using the COBYLA optimiser. The cost function
$C_{t}(\vec{\theta})=\text{Tr}(\rho^{2})-\text{Tr}(\rho_{t}(\vec{\theta})^{2}).$ (12)
at each step $t$ is calculated for the ansatz which outputs a state $\rho_{t}(\vec{\theta})$.
Appendix C Analysis of the Haar random input quantum states
In this section, we analyse the 2-qubit quantum states sampled from random_density_matrix of IBM qiskit’s quantum_info module. Here, we sample $100000$ Haar random quantum states and observe how the eigenvalues and the qubits’ conditional entropy change with the state’s entanglement.
Figure 7: The difference in conditional entropy converges towards the zero as the entanglement of the sampled input state increases. Meanwhile, as the input state gets closer to maximally entangled state (with concurrence $1$), the magnitude of the largest eigenvalue increases while the other eigenvalue decreases.
We see that the difference in conditional entropy in qubits of the input state varies evenly on either side of $S_{AB}^{\rho_{AB}}=0$ where $A$ and $B$ are subsystems containing one qubit each and $\rho_{AB}$ is the sampled Haar random state. Meanwhile, the area of the distribution of $S_{AB}^{\rho_{AB}}$ on the ether side of zero entropy converges towards zero as the input state entanglement increases. Furthermore, the magnitude of the input quantum state’s largest eigenvalue increases with the increase in entanglement, and all the other eigenvalues converge towards 00.

### v2

@@
 RL-QAS involves optimising QAS using reinforcement learning techniques.
-This study investigates RL-QAS for crafting ansatzes tailored to the variational quantum state diagonalization problem.
-Our investigation includes a comprehensive analysis of various dimensions, such as the entanglement thresholds of the resultant states, the impact of initial conditions on the performance of RL-agent, the phase change behavior of correlation in concurrence bounds, and the discrete contributions of qubits in deducing eigenvalues through conditional entropy metrics.
-We leverage these insights to devise an optimal, admissible QAS to diagonalize random quantum states.
+This study investigates RL-QAS for crafting ansatzes tailored to the variational quantum state diagonalisation problem.
+Our investigation includes a comprehensive analysis of various dimensions, such as the entanglement thresholds of the resultant states, the impact of initial conditions on the performance of RL-agent, the phase change behaviour of correlation in concurrence bounds, and the discrete contributions of qubits in deducing eigenvalues through conditional entropy metrics.
+We leverage these insights to devise entanglement-guided admissible ansatz in QAS to diagonalise random quantum states using optimal resources.
 Furthermore, the methodologies presented herein offer a generalised framework for constructing reward functions within RL-QAS applicable to variational quantum algorithms.
@@
 (2021).
-QAS can also used to determine non-parametric circuits as an approach for quantum program synthesis Sarkar (2022).
+QAS can also be used to determine non-parametric circuits as an approach for quantum program synthesis Sarkar (2022).
 In this research, however, we will focus on QAS in the context of VQA.
@@
 Given the vast potential configuration space of quantum circuits, the process of identifying a suitable ansatz is a challenging problem.
-In previous works such as Ostaszewski et al.
+In Ostaszewski et al.
 (2021); Kundu et al.
@@
 (2022), the search for the quantum circuits that solve various VQA has been automated using reinforcement learning (RL) techniques Sutton and Barto (2018).
-We particularly call quantum ansatzes capable of expressing the solution of VQA an admissible ansatz.
+We particularly call ansatzes capable of solving a VQA an admissible ansatz.
 The automated search for such admissible ansatz using the RL-agent is termed reinforcement learning-assisted quantum architecture search (RL-QAS).
@@
 One of the ways of providing this information is by engineering a reward function.
-In RL, the formulation of the reward function and its engineering can be sparse and dense.
+In RL, the formulation of the reward function and its engineering can be sparse or dense.
 The choice of optimal reward function depends on trial and error by investigating the performance of the agent under various reward configurations.
@@
 (2024).
-In the VQSD algorithm, the problem-inspired ansatz is based on density matrix exponentiation Lloyd et al.
+In the VQSD algorithm LaRose et al.
+(2019), the problem-inspired ansatz is based on density matrix exponentiation Lloyd et al.
 (2014), but its depth and number of gates increase exponentially with the number of qubits.
@@
 The RL-agent proposes $E_{s}<E_{tot}$ admissible ansatzes that solves the VQSD problem.
-Our investigation shows that the concurrence of the $s$ ansatzes lies between an upper and lower bound, independent of the different initialisation of the double deep-Q network (DDQN), where each initialisation of DDQN corresponds to a different random initialisation of the weights.
+Our investigation shows that the concurrence of the $E_{s}$ ansatz lies between an upper and lower bound, independent of the different initialisation of the double deep-Q network (DDQN), where each initialisation of DDQN corresponds to a different random initialisation of the weights.
 Moreover, we show that the upper and lower bounds are anti-correlated with respect to the increment in the initial state entanglement.
@@
 Further, our investigation reveals that the RL-agent can generate high concurrence admissible ansatz with fewer 2-qubit gates and circuit depth compared to ansatzes with very low concurrence.
+This is used as a guiding principle to design a two-part ansatz, where an entanglement enhancing initial block adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$.
 Furthermore, we quantify the contribution of individual qubits of the RL-ansatz using conditional quantum entropy.
@@
   1. 1.
-    Investigating admissible ansatzes of reinforcement learning based quantum architecture search problem from a quantum information theory perspective.
-We focus on the RL-VQSD algorithm, where the proposed ansatzes diagonalise random quantum states.
+    Demonstrating that the entanglement of the states generated by the RL-agent-proposed admissible ansatzes lies within a lower and upper concurrence bound.
+The upper and lower bounds of concurrence remain consistent across various weight initialisation of the DDQN.
+Furthermore, we show that these bounds are anti-correlated with respect to increasing entanglement until concurrence surpasses $0.322$; after this point, the anti-correlation turns into correlation.
+Hence, we introduce the concurrence $0.322$ as the phase change point.
   2. 2.
-    Demonstrating that the entanglement of the states generated by the RL-agent-proposed admissible ansatzes lies within a lower and upper concurrence bound.
+    Investigation of ansatzes in upper and lower bounds reveals that the optimal ansatz for VQSD generates strongly entangled states.
+The RL-agent requires fewer gates and circuit depth to generate such states to produce strongly entangled admissible ansatzes than weakly entangled ansatzes.
+We then utilise this information to minimise RL-agent and environment (which is the configuration of ansatz) interaction, thereby optimising the convergence time of the RL-VQSD.
+Specifically, we propose a two-part ansatz, where an entanglement enhancing initial block adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$.
   3. 3.
-    The upper and lower bounds of concurrence remain consistent across various weight initialisation of the DDQN.
+    We evaluate individual qubits’ contribution in diagonalising 2-qubit quantum states using conditional quantum entropy.
+Our investigation shows that the qubits have equal contributions on average in diagonalising random quantum states.
   4. 4.
-    Bounds are anti-correlated with respect to increasing entanglement until concurrence surpasses $0.322$; after this point, the anti-correlation turns into correlation.
-Hence, we introduce the concurrence $0.322$ as the phase change point.
-  5. 5.
-    Investigation of ansatzes in upper and lower bounds, showing optimal ansatz for VQSD generates high concurrence states.
-  6. 6.
-    RL-agent requires fewer gates and circuit depth to produce highly entangled admissible ansatzes than weakly entangled ansatzes.
-We utilise this information to minimise RL-agent and environment (which is the configuration of ansatz) interaction, thereby optimising the convergence time of the RL-VQSD.
-  7. 7.
-    We evaluate individual qubits’ contribution in diagonalising 2-qubit quantum states using conditional quantum entropy.
-  8. 8.
-    Our investigation shows that the qubits have equal contributions on average in diagonalising random quantum states.
-  9. 9.
     We further investigate the correlation in conditional entropy between qubits for different inferred eigenvalues.
 It requires mild anti-correlation for the first two largest eigenvalues and strong anti-correlation for the smallest eigenvalues.
-  10. 10.
-    The majority of admissible ansatzes lie in a mild anti-correlation regime, explaining VQSD’s limitation in finding the smallest eigenvalues.
+Most admissible ansatzes lie in a mild anti-correlation regime, explaining VQSD’s limitation in finding the smallest eigenvalues.
 III Main results
@@
 III.1.1 DDQN weight independent entanglement bound
-To benchmark the performance of the RL agent for a specific problem, it is ideal to take multiple initialisation of the DDQN (see Appendix B for a brief description).
+To benchmark the performance of the RL-agent for a specific problem, it is ideal to take multiple initialisations of the DDQN (see Appendix B for a brief description).
 state max. concurrence min. concurrence
@@
 9 0.831 0.011 0.302 0.010
-Table 1: In this table, we present the average and standard deviation of the concurrence of the final state after evolution through the circuit for five different initialisations of the DDQN wrights for different quantum seed states.
-To efficiently investigate the upper and lower bounds of concurrence Hill and Wootters (1997); Wootters (1998) produced by the RL ansatz, we first sample different random quantum states, then diagonalize them using RL-VQSD Kundu et al.
-(2024) for five different initialisation of DDQN weights.
-Thereafter, the agent suggests different possible configurations for the RL-ansatz, which diagonalizes the sampled random quantum states.
+Table 1: The average and standard deviation of the final state’s maximum (max.
+concurrence) and minimum concurrence (min.
+concurrence) after evolution through the circuit for five different initialisations of the DDQN weights for different quantum seed states.
+To obtain the results, we run RL-VQSD to diagonalize $2$-qubit random quantum states for $10000$ episodes.
+The agent then proposes $E_{s}<10000$ admissible ansatzes, which generate $E_{s}$ states after the random quantum state evolves them.
+The max.
+(min.) concurrence is attributed to the ansatz that generates a state with the highest (lowest) concurrence following the evolution of the random quantum state.
+To efficiently investigate the upper and lower bounds of concurrence Hill and Wootters (1997); Wootters (1998) of the admissible ansatz, we calculate the concurrence of admissible ansatz, which is defined by the entanglement induced by the admissible ansatz when initialised with |$\ket{00}$.
+We first sample different random quantum states, then diagonalise them using RL-VQSD Kundu et al.
+(2024) for five different initialisations of DDQN weights.
+Thereafter, the agent suggests different possible configurations for the RL-ansatz, which diagonalises the sampled random quantum states.
 For a specific state, we collect such valid quantum circuits and find the concurrence of the state obtained after evolving through the RL-ansatz.
 It should be noted that finding the concurrence of the evolved state indicates how much the RL-ansatz induces entanglement to the input sampled state.
-See Appendix C for further analysis of the sampled random quantum states.
+See Appendix C for further analysis of the sampled random quantum states, where we show that in qiskit, random_density_matrix even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low.
 The results are presented in Table 1.
@@
 III.1.2 Anti-correlation between entanglement bounds
-Figure 1: Reciprocal behaviour in the upper and lower bounds of concurrence with respect to the increasing concurrence in the input random state.
-A clear anti-correlation between entanglement minima and maxima which breaks down after $0.3$.
-From the interpolating function, the Pearson correlation coefficient (PCC) for the first eleven points is -0.982, while that for the remaining points is 0.049.
-At this point, the correlation after the first eleven points changes from -0.055 to 0.049.
-There is a phase transition from weak anti-correlation to weak correlation.
 Corresponding to a predefined threshold, the agent proposes different architectures of quantum circuits that diagonalise the quantum state.
@@
 By sorting out the quantum circuit with maximum and minimum entanglement, we observe their variation with the entanglement of the input state.
-It has been proved in the previous section that the upper and lower bounds of entanglement are independent of different initialisation of DDQNs; hence, we for further benhamarking we can constrain ourselves to just one initialisation of the DDQN weights.
+It has been proved in the previous section that the upper and lower bounds of entanglement are independent of different initialisation of DDQNs; hence, for further benchmarking, we can constrain ourselves to just one initialisation of the DDQN weights.
+Figure 1: Reciprocal behaviour in the upper and lower bounds of concurrence with respect to the increasing concurrence in the input random state.
+For the first eleven points, the Pearson correlation coefficient (PCC) is $-0.982$, denoting a strong anti-correlation between the upper and lower bounds of concurrence.
+The strong anti-correlation turns into a mild correlation after concurrence of the initial state surpasses $0.322$.
+The correlation after the first eleven points changes from $-0.055$ to $0.049$.
+There is a phase transition from weak anti-correlation to weak correlation.
 In Figure 1, we illustrate the variation of the entanglement bounds with the entanglement of the input quantum state.
@@
 See Appendix A for a brief description.
-We see that the upper and lower bounds of entanglement are initially strongly anti-correlation when the input state entanglement is below $0.3$ and then strongly correlate afterwards, indicating a phase change in the upper and lower bounds of entanglement.
-Figure 2: The PCC indicating Pearson correlation coefficient is calculated over an interval $i<k<j$.
-The coefficient is described as in Equation 1, where $i$ is minimum, $j$ is maximum concurrence, and $k$ splits the intermediate values of concurrence into two parts.
-Hence, $i$ and $j$ are fixed.
-The $w_{ik}\in[0,1]$ is the weight of the first interval which defines how much each of the two intervals contributes when varying $k$.
-We vary the cardinality of the two intervals, causing variation in the weights.
-In the figure, we consider $i=0.119$ and $j=0.402$ and when $k=0.322$, i.e.
-$i<x\leq 0.322$, the weight corresponds to the first interval.
+We see that the upper and lower bounds of entanglement are initially strongly anti-correlation when the input state entanglement is below $0.322$ and then strongly correlate afterwards, indicating a phase change in the upper and lower bounds of entanglement.
+Figure 2: We present the phase transition of correlation between the upper and lower bounds of concurrence for the admissible RL-VQSD ansatz.
+To show this, we split the total range of the concurrence of the input state in two intervals $\mathbbm{u}_{1}\coloneqq[i,k)$ and $\mathbbm{u}_{2}\coloneqq[k,j]$, where $k$ is a free parameter and plot the variation of the Pearson correlation coefficient between the upper and lower bound on the concurrence of the ansatz corresponding to these two intervals.
+We denote by $\textrm{PCC}_{ik}$ and $\textrm{PCC}_{kj}$ the correlation coefficients in the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$ respectively.
+The phase transition point is observed at $k=0.322$ where the concurrence bounds change from mild anti-correlation to mild correlation for the interval $\mathbbm{u}_{2}$ while remaining strongly anti-correlated in the interval $\mathbbm{u}_{1}$.
 Next, we obtain the variation of PCC between the upper and lower bound on the concurrence of the initial state.
@@
 For both the ranges $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$, we introduce the following correlation function:
-$\eta^{k}_{w_{ik}}=w_{ik}\textrm{PCC}^{ik}_{ul}+(1-w_{ik})\textrm{PCC}^{kj}_{ul},$ (1)
-where $w_{ik}\in\{0,1\}$ and $\textrm{PCC}^{ik}_{ul}$, $\textrm{PCC}^{kj}_{ul}$ are the value of PCC between the upper and lower bounds of concurrence of the ansatz for the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$ respectively.
+$\eta^{k}_{w_{ik}}=w_{ik}\textrm{PCC}_{ik}+(1-w_{ik})\textrm{PCC}_{kj},$ (1)
+where $w_{ik}\in\{0,1\}$ and $\textrm{PCC}_{ik}$, $\textrm{PCC}_{kj}$ are the value of PCC between the upper and lower bounds of concurrence of the ansatz for the intervals $\mathbbm{u}_{1}$ and $\mathbbm{u}_{2}$ respectively.
+For the sake of simplicity we denote $\textrm{PCC}_{ik}$ as $\textrm{PCC}_{ik}^{ul}$ and also $\textrm{PCC}_{kj}$ as $\textrm{PCC}_{kj}^{ul}$.
 Note 1.
-[…]
+We observe that depending on the values of $w_{ik}$, $\eta^{k}_{w_{ik}}$ can take values either $\textrm{PCC}_{ik}$ or $\textrm{PCC}_{kj}$.
+For $\eta^{k}_{w_{ik}}=\textrm{PCC}_{ik}$, we obtain the amount of (anti-)correlation between the lower and upper bounds of concurrence of the ansatz for the interval $\mathbbm{u}_{1}$.
+While for $\eta^{k}_{w_{ik}}=\textrm{PCC}_{kj}$, we obtain the amount of (anti-)correlation for the interval $\mathbbm{u}_{2}$.
+We leave proving $\eta^{k}_{w_{ik}}$ is a valid measure of correlation for arbitrary values of $w_{ik}\in(0,1)$ as an open problem.
 We present in Figure 2, the variation of $\eta_{0}^{k}$ (in blue) and $\eta_{1}^{k}$ (in red) for different values of $k$.
@@
 This indicates strong cumulative anti-correlation between the upper and lower bounds of concurrence.
-For $\eta_{0}^{k}$, we observe a gradual change of $\text{PCC}^{kj}_{ul}$ from -0.582 to 1.
+For $\eta_{0}^{k}$, we observe a gradual change of $\text{PCC}_{kj}$ from -0.582 to 1.
 We note that $\eta_{0}^{k}$ changes from $-0.0451$ to $0.279$ for $k\in[0.313,0.379]$ with $\eta_{0}^{k}=0$ at $k=k_{\ast}=0.322$.
@@
 While for values of $k>k_{\ast}$, there is (anti-)correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $(\mathbbm{u}_{2})~{}\mathbbm{u}_{1}$.
-III.1.3 Agent and environment interaction reduction based on entanglement bound
+III.1.3 Enhancing the performance of agent based on initial state entanglement
 Here, we provide a detailed investigation of the admissible ansatzes produced by the RL-agent in terms of ansatz depth, the total number of 2-qubit gates and the total number of gates in the upper and lower bounds of concurrence.
-Figure 3: We observe that the RL-agent can generate ansatz with very high concurrence with less number of 2-qubit gates and circuit depth compared to the ansatzes with very low concurrence.
+Figure 3: The RL-agent can generate an ansatz with strong entanglement with fewer 2-qubit gates and circuit depth compared to the ansatzes with weak entanglement.
 As the main aim of RL-VQSD is to diagonalise unitary with a small number of gates and depth, our analysis shows that if we particularly focus on admissible ansatzes near the upper bound of concurrence, we can not only reduce the computational cost of RL-VQSD severely but find optimal admissible ansatzes.
@@
 This observation is crucial in reducing the subspace of all possible admissible ansatzes.
-The main motivation of RL-QAS in VQA is to find ansatzes that solve a VQA with a minimum number of gates and circuit depth; in the case of RL-VQSD, we can narrow down the search for an optimal admissible ansatz near the upper bound of entanglement by encouraging the RL-anent to produce ansatzes with very high concurrence and penalise for the other structures of ansatz.
+The main motivation of RL-QAS in VQAs is to find ansatzes that solve a VQA with a minimum number of gates and circuit depth; in the case of RL-VQSD, this observation can help us to narrow down the search for an optimal admissible ansatz in the two following ways.
+a. Entanglement dependent initialisation of RL-state
+In RL-VQSD Kundu et al.
+(2024), the RL-state that encodes the ansatz initiates from an empty circuit (i.e., in circuit 2 $U_{\text{RL-agent}}=\mathds{1}$).
+We showed that the optimal admissible ansatz, on average, lies in the upper bound of concurrence, motivating us to initialise the quantum state (that is diagonalised using RL-VQSD) closest to the upper bound of entanglement.
+To do this, the ansatz is divided into two parts
+[…]
+where the first, Entanglement Enhancing (EE) block, helps the diagonalising state to achieve the highest entanglement before it is fed to the second part, $U_{\text{RL-agent}}$, which is decided by the RL-agent.
+For example, for seed $27$ (see Appendix C for details), we get a random quantum state with concurrence $0.198$, and using the EE, we can increase the input state concurrence to $0.215$ at $\theta=0.5$ as shown in Fig.
+4.
+Figure 4: The concurrence of input state with $\theta$ of CRX($\theta$) gate in Entanglement Enhancing increases the concurrence of the input random state from $0.198$ to $0.215$ when $\theta=0.5$.
+Our investigation reveals that turning on the EE block with $\theta=0.5$ enhances the performance of the RL-agent by $2$-times compared to the default setting at $\theta=0$.
+The results are summarised in the Tab.
+2 where we see that the number of successful episodes and the reward acquired (where the reward function is in the form defined in Eq.
+12) by the agent in the RL-VQSD process increase by $2$-fold while the EE block is activated for $\theta=0.5$ compared to default settings at $\theta=0$.
+Table 2: Turning on the Entanglement Enhancing (EE) mode (denoted as Enhanced ($\theta=0.5$)) enhances the learning of the performance of the RL-agent in the absence of the EE mode (denoted as Default ($\theta=0$)).
+In the first $3500$, the Enhanced ($\theta=0.5$) collects $2.08$ times more reward (tot.
+rwd.
+coll.) and $2.22$ times more successful episodes (Tot.
+succ.
+ep.) than Default ($\theta=0$).
+Moreover, the average number of 2-qubit (2q gate) and depth over the first $100$ successful episodes reduced in the case of Enhanced ($\theta=0.5$) settings compared to Default ($\theta=0$), showing enhanced learning of the RL-agent for entanglement enhanced input state.
+EE mode Tot. succ. Tot. rwd. 1st 100 succ. ep. (avg.)
+($\theta$) ep. coll. 1q gate 2q gate depth
+Enhanced ($\theta=0.5$) $858$ $4.45\times 10^{5}$ $15.41$ $4.02$ $14.6$
+Default ($\theta=0$) $387$ $2.14\times 10^{5}$ $14.71$ $5.29$ $15.59$
+Meanwhile, in the Tab. 3, we show that the EE block also helps to optimise the cost function and the minimum number of 2-qubit gates in the optimal admissible ansatz compared to the default settings.
+Table 3: Turning on the Entanglement Enhancing (EE) mode (denoted as Enhanced ($\theta=0.5$)) optimises the number of minimum number of 2-qubit gates (Min.
+2q gate) in RL-ansatz better than in the absence of the EE mode (denoted as Default ($\theta=0$)).
+Moreover, the Enhanced ($\theta=0.5$) mode minimises the cost function better than the Default ($\theta=0$) mode.
+EE mode ($\theta$) Cost function Min. 1q gate Min. 2q gate
+Enhanced ($\theta=0.5$) $5.12\times 10^{-7}$ $14$ $2$
+Default ($\theta=0$) $6.08\times 10^{-7}$ $11$ $5$
+The aforementioned observations lead us to the conclusion that maximising the entanglement of the input state through the initialisation strategy of Fig.
+2 enhances the learning process of the RL-agent and facilitates the further optimisation of the admissible ansatzes compared to the start-of-art Kundu (2024).
 III.2 Conditional quantum entropy
@@
 For a detailed discussion and notation, see Appendix A.
-Figure 4: Relative change in the contribution of qubits 0 and 1 as a function of change in entanglement. Weight of the positive part $5.524$ and that of the negative part is $-4.854$.
+Figure 5: Relative change in the contribution of qubits 0 and 1 as a function of change in entanglement.
+We quantify the relative contribution of individual qubits to the RL-VQSD ansatz using $\Delta_{q_{0}}^{q_{1}}$ (see Eq.
+(3)).
+For the change in concurrence due to the ansatz, the distribution of $\Delta_{q_{0}}^{q_{1}}$ is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$.
+The magnitude of the cumulative weight of all data points in the region $\Delta_{q_{0}}^{q_{1}}<0$ is $4.854$ while that for data points in the region $\Delta_{q_{0}}^{q_{1}}>0$ is $5.524$.
+This indicates equal contribution from both the qubits to the RL-VQSD ansatz for the range of change in entanglement.
 Let the change in the conditional entropy of the qubit $q_{0}(q_{1})$ after evolving through the VQSD ansatz be given by $\Delta S^{\rho\to{\rho^{\prime}}}_{q_{0}|q_{1}}(\Delta S^{\rho\to{\rho^{\prime}}}_{q_{1}|q_{0}})$.
@@
 We observe that the distribution of $\Delta_{q_{0}}^{q_{1}}$ is bounded in the interval $[-0.7,0.7]$ and is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$.
-This implies an equal contribution to the VQSD ansatz from both the qubits for the range of change in entanglement.
+The magnitude of the cumulative weight of all data points in the region $\Delta_{q_{0}}^{q_{1}}<0$ is $4.854$ while that for data points in the region $\Delta_{q_{0}}^{q_{1}}>0$ is $5.524$.
+This indicates equal contribution from both the qubits to the RL-VQSD ansatz for the range of change in entanglement.
 In the region of $\Delta_{q_{0}}^{q_{1}}>0$, the relative contribution of $q_{0}$ first increases and then decreases.
@@
 Also, for $\Delta_{c}<0.2$, $\Delta_{q_{0}}^{q_{1}}$ increases monotonically with increase in $\Delta_{c}$.
-III.2.2 Problem with variational quantum state diagonalization
+III.2.2 Problem with variational quantum state diagonalisation
 In this section, we investigate the contribution of individual qubits to the performance of the RL-VQSD algorithm.
@@
 Then, as the magnitude of the smallest eigenvalue increases, the strong anti-correlation turns into mild anti-correlation.
-Figure 5: The conditional entropy correlation between the first and the second qubit decreases as the magnitude of the largest and the 2nd largest eigenvalues increases.
-Meanwhile, the same correlation between the qubits increases with the magnitude of the 3rd and the smallest eigenvalues.
+Figure 6: The conditional entropy correlation between the first and the second qubit decreases as the magnitude of the largest and the second largest eigenvalues increases.
+Meanwhile, the same correlation between the qubits increases with the magnitude of the third and the smallest eigenvalues.
 The RL-ansatz proposes specific structures of admissible ansatzes, and for each such ansatz, it is not possible to have two different kinds of correlations.
 Hence, our investigation suggests that it is not feasible to find the smallest eigenvalues with the same ansatz that provides a good approximation of the largest eigenvalues.
-The correlation of conditional entropy between qubits is calculated using Pearson correlation function (PCC) and is the median over all the configuration of RL-ansatz proposed by the agent.
+The correlation of conditional entropy between qubits is calculated using Pearson correlation coefficient (PCC) and is the median over all the configuration of RL-ansatz proposed by the agent.
 The $\bigstar$ signifies the eigenvalues corresponding to the same state.
@@
 This observation, along with the observation in Figure 6 certify that RL-agent only proposes a particular configuration of admissible ansatz that can find only the largest eigenvalues.
-The average correlation among all the $21$ states is $-0.4263$, which is in the regime of mild anti-correlation.
+The average correlation among all the $21$ states is $-0.4263$ (horizontal black line), which is in the regime of mild anti-correlation.
 IV Discussion
@@
 We effectively utilise these observations to greatly reduce the RL-agent and environment interaction, i.e., the computational time needed to solve the RL-VQSD problem.
-Additionally, the admissible ansatzes not only diagonalize the random quantum states with high accuracy but can be used to generate close to maximum entangled states from the vacuum state.
+Specifically, we propose an entanglement enhancing initial block that adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$ Additionally, the admissible ansatzes not only diagonalise the random quantum states with high accuracy but can be used to generate close to maximum entangled states from the vacuum state.
 Furthermore, we quantify the contribution of each qubit in the admissible ansatz using conditional quantum entropy.
@@
 From the observation, most of the admissible ansatz has mild anti-correlation among the qubits, which explains why it is easier for the diagonalisation algorithm to find the largest eigenvalues than the smallest ones.
-For future work, it would be interesting to see the implication of introducing entanglement Horodecki et al.
+For future work, it would be interesting to conduct more comprehensive study of the implication of introducing entanglement Horodecki et al.
 (2009) or nonlocality Brunner et al.
-(2014); Sadhu and Das (2023)-based reward functions of the RL-agent for solving the VQSD problem.
+(2014); Sadhu and Das (2023)-based reward functions of the RL-agent in QAS.
 Specifically, given that most of the admissible ansatzes have high values of concurrence, it is expected that an entanglement or non-locality-based reward function may further reduce the time required to find the optimal configurations.
@@
 A.
-Sadhu acknowledges the PhD fellowship from the Raman Research Institute, Bangalore.
-AK would like to thank Mateusz Ostaszewski for the valuable discussions.
+Sadhu acknowledges the Ph.D.
+fellowship from the Raman Research Institute, Bangalore.
+AK would like to thank Jarosław Miszczak and Mateusz Ostaszewski for their valuable comments on the manuscript and discussions.
 This research was partially supported by PL-Grid Infrastructure Grant Number PLG/2023/016205.
@@
 We quantify the amount of correlation between two random variables $r_{1}$ and $r_{2}$ using the Pearson correlation coefficient (PCCLee Rodgers and Nicewander (1988) defined as:
-$\textrm{PCC}_{r_{1}r_{2}}=\frac{\mathrm{E}[(r_{1}-\mu_{r_{1}})(r_{2}-\mu_{r_{2}})]}{\sigma_{r_{1}}\sigma_{r_{2}}},$ (9)
+$\textrm{PCC}^{r_{1}r_{2}}=\frac{\mathrm{E}[(r_{1}-\mu_{r_{1}})(r_{2}-\mu_{r_{2}})]}{\sigma_{r_{1}}\sigma_{r_{2}}},$ (10)
 where $\mu_{r_{j}}$, $\sigma_{r_{j}}$ are the mean and standard deviation of the distribution $r_{j}$ with $j\in\{1,2\}$ and $\mathrm{E}[.]$ is an operator that returns the expectation value of a random variable.
-Note that the Pearson correlation coefficient (PCC) is symmetric: $\textrm{PCC}_{r_{1}r_{2}}=\textrm{PCC}_{r_{2}r_{1}}$.
+Note that the Pearson correlation coefficient (PCC) is symmetric: $\textrm{PCC}^{r_{1}r_{2}}=\textrm{PCC}^{r_{2}r_{1}}$.
 The value of PCC ranges in between $-1$ to $+1$.
@@
 $\rho^{\prime}=U(\vec{\theta}_{\text{opt}})\rho U(\vec{\theta}_{\text{opt}})^{\dagger}=\rho_{\text{diag}},$
-where $\rho_{\text{diag}}$ is $\rho$ diagonalized in its eigenbasis and $\vec{\theta}_{\text{opt}}$ are the optimal angles.
-Classical gradient-based methods such as SPSA and Gradient Descent, or gradient-free optimisation Powell (1994, 2006), can be utilized in the training process.
+where $\rho_{\text{diag}}$ is $\rho$ diagonalised in its eigenbasis and $\vec{\theta}_{\text{opt}}$ are the optimal angles.
+Classical gradient-based methods such as SPSA and Gradient Descent, or gradient-free optimisation Powell (1994, 2006), can be utilised in the training process.
 (2) In the eigenvalue readout subroutine, using the optimised unitary $U(\vec{\theta}_{\text{opt}})$ and one copy of state $\rho$, one can extract all the eigenvalues for low-rank states or the largest eigenvalues for full-rank states.
@@
 Figure 8: The difference in conditional entropy converges towards the zero as the entanglement of the sampled input state increases.
-Meanwhile, as the input state gets closer to maximally entangled state (with concurrence $1$), the magnitude of the largest eigenvalue increases while the other eigenvalue decreases.
-We see that the difference in conditional entropy in qubits of the input state varies evenly on either side of $S_{AB}^{\rho_{AB}}=0$ where $A$ and $B$ are subsystems containing one qubit each and $\rho_{AB}$ is the sampled Haar random state.
+Meanwhile, as the input state gets closer to the maximally entangled state (with concurrence $1$), the magnitude of the largest eigenvalue increases while the other eigenvalue decreases.
+In Fig.
+8, see that the difference in conditional entropy in qubits of the input state varies evenly on either side of $S_{AB}^{\rho_{AB}}=0$ where $A$ and $B$ are subsystems containing one qubit each and $\rho_{AB}$ is the sampled Haar random state.
 Meanwhile, the area of the distribution of $S_{AB}^{\rho_{AB}}$ on the ether side of zero entropy converges towards zero as the input state entanglement increases.
 Furthermore, the magnitude of the input quantum state’s largest eigenvalue increases with the increase in entanglement, and all the other eigenvalues converge towards 00.
+It also shows that even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low, justifying the low density of points after $0.6$ in Fig.
+1.
+Sampling a random quantum state
+If $n$ is the number of qubits and $s$ is the random seed to sample, then we can sample a state from random_density_matrix of IBM qiskit’s quantum_info module (in python) as follows
+[…]
+The state used in III.1, is sampled using $n=2$ and $s=27$

### v3

@@
 These algorithms demonstrate significant potential for realising quantum computational advantage.
-A fundamental aspect of VQAs involves formulating expressive and efficient quantum circuits (namely ansatz) and automating the search of such ansatz is known as quantum architecture search (QAS).
-RL-QAS involves optimising QAS using reinforcement learning techniques.
-This study investigates RL-QAS for crafting ansatzes tailored to the variational quantum state diagonalisation problem.
-Our investigation includes a comprehensive analysis of various dimensions, such as the entanglement thresholds of the resultant states, the impact of initial conditions on the performance of RL-agent, the phase change behaviour of correlation in concurrence bounds, and the discrete contributions of qubits in deducing eigenvalues through conditional entropy metrics.
-We leverage these insights to devise entanglement-guided admissible ansatz in QAS to diagonalise random quantum states using optimal resources.
+A fundamental aspect of VQAs involves formulating expressive and efficient quantum circuits (namely ansatz), and automating the search of such ansatz is known as quantum architecture search (QAS).
+Recently, reinforcement learning (RL) techniques is utilized to automate the search for ansatzes, known as RL-QAS.
+This study investigates RL-QAS for crafting ansatz tailored to the variational quantum state diagonalisation problem.
+Our investigation includes a comprehensive analysis of various dimensions, such as the entanglement thresholds of the resultant states, the impact of initial conditions on the performance of RL-agent, the phase transition behaviour of correlation in concurrence bounds, and the discrete contributions of qubits in deducing eigenvalues through conditional entropy metrics.
+We leverage these insights to devise an entanglement-guided admissible ansatz in QAS to diagonalise random quantum states using optimal resources.
 Furthermore, the methodologies presented herein offer a generalised framework for constructing reward functions within RL-QAS applicable to variational quantum algorithms.
@@
 We particularly call ansatzes capable of solving a VQA an admissible ansatz.
-The automated search for such admissible ansatz using the RL-agent is termed reinforcement learning-assisted quantum architecture search (RL-QAS).
+The automated search for such an admissible ansatz using the RL-agent is termed reinforcement learning-assisted quantum architecture search (RL-QAS).
 In this setting, after a certain number of episodes, the RL agent returns possible configurations of quantum ansatz that solve the problem.
@@
 The choice of optimal reward function depends on trial and error by investigating the performance of the agent under various reward configurations.
-In this article, we investigate the admissible ansatzes proposed by the RL-agent from the perspective of quantum information theory Wilde (2013).
+In this article, we numerically investigate the admissible ansatzes proposed by the RL-agent from the perspective of quantum information theory Wilde (2013).
 For the investigation, we specifically focus on the reinforcement learning enhanced variational quantum state diagonalisation (RL-VQSD) problem Kundu et al.
@@
 In RL-VQSD, the search for the optimal configuration of ansatz is done by using a double deep-Q network (DDQN).
-See Appendix B for further details on the VQSD algorithm, the agent and environment specifications.
+See Section III for further details on the VQSD algorithm, the agent and environment specifications.
 The RL-agent proposes $E_{s}<E_{tot}$ admissible ansatzes that solves the VQSD problem.
+The primary motivation of our work is to filter potential admissible ansatzes based on concurrence, using the highest and lowest entanglement bounds.
+This helps identify the range of optimal ansatzes and determine if increasing entanglement can improve the agent’s performance.
+Another goal is to see if we can find an admissible ansatz that both diagonalizes and enhances the entanglement of the vacuum state.
 Our investigation shows that the concurrence of the $E_{s}$ ansatz lies between an upper and lower bound, independent of the different initialisation of the double deep-Q network (DDQN), where each initialisation of DDQN corresponds to a different random initialisation of the weights.
-Moreover, we show that the upper and lower bounds are anti-correlated with respect to the increment in the initial state entanglement.
-The anti-correlation turns into a mild, and later strong, correlation when the initial state entanglement goes beyond $0.322$, indicating a phase change in correlation between the upper and lower bounds of the concurrence.
+Moreover, through the numerical analysis we show that the upper and lower bounds are anti-correlated with respect to the increment in the initial state entanglement.
+The anti-correlation turns into a mild, and later strong, correlation when the initial state entanglement goes beyond $0.322$, indicating a phase transition in correlation between the upper and lower bounds of the concurrence.
 The phase transition provides us with an in-depth insight into the relation between the input and the output of the VQSD algorithm.
-Further, our investigation reveals that the RL-agent can generate high concurrence admissible ansatz with fewer 2-qubit gates and circuit depth compared to ansatzes with very low concurrence.
+Further, our investigation reveals that the RL-agent can generate high concurrence admissible ansatzes with fewer 2-qubit gates and circuit depth compared to ansatzes with very low concurrence.
 This is used as a guiding principle to design a two-part ansatz, where an entanglement enhancing initial block adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$.
@@
 In Sec.
-III.1, we analyse the upper and lower bounds on the concurrence of the admissible ansatzes that solve the VQSD problem.
+III, we present an overview of the reinforcement learning enhanced variational quantum state diagonalisation algorithm.
+We present an analysis of the sampled 2-qubit Haar-random quantum states in Sec.
+IV.
+We present the main results in Sec.
+V.
 Specifically, in Sec.
-III.1.1 we present the variation of the entanglement bounds for different weight initialisation of the DDQN.
+V.1, we analyse the upper and lower bounds on the concurrence of the admissible ansatzes that solve the VQSD problem.
+In Sec.
+V.1.1 we present the variation of the entanglement bounds for different weight initialisation of the DDQN.
 We observe the correlation properties of the entanglement bounds dependent on the concurrence of the initial state in Sec.
-III.1.2.
+V.1.2.
 In Sec.
-III.1.3, we present the dependence of the entanglement bounds on the number of gates and circuit depth of the ansatz.
+V.1.3, we present the dependence of the entanglement bounds on the number of gates and circuit depth of the ansatz.
 The analysis of the contribution of individual qubits to the admissible ansatz is provided in Sec.
-III.2.
+V.2.
 In particular, we present the contribution of each qubit for change in the entanglement of the ansatz in Sec.
-III.2.1.
+V.2.1.
 In Sec.
-III.2.2, we explain why RL-VQSD can efficiently find the largest eigenvalues of the initial state but not the smallest eigenvalues.
+V.2.2, we explain why RL-VQSD can efficiently find the largest eigenvalues of the initial state but not the smallest eigenvalues.
 We provide concluding remarks and discuss possible future directions in Sec.
-IV.
+VI.
 II Contributions
@@
   1. 1.
-    Demonstrating that the entanglement of the states generated by the RL-agent-proposed admissible ansatzes lies within a lower and upper concurrence bound.
+    We numerically demonstrate that the entanglement of the states generated by the RL-agent-proposed admissible ansatzes lies within a lower and upper concurrence bound.
 The upper and lower bounds of concurrence remain consistent across various weight initialisation of the DDQN.
 Furthermore, we show that these bounds are anti-correlated with respect to increasing entanglement until concurrence surpasses $0.322$; after this point, the anti-correlation turns into correlation.
-Hence, we introduce the concurrence $0.322$ as the phase change point.
+Hence, we introduce the concurrence $0.322$ as the phase transition point.
   2. 2.
@@
   3. 3.
-    We evaluate individual qubits’ contribution in diagonalising 2-qubit quantum states using conditional quantum entropy.
+    We numerically evaluate individual qubits’ contribution in diagonalising 2-qubit quantum states using conditional quantum entropy.
 Our investigation shows that the qubits have equal contributions on average in diagonalising random quantum states.
@@
     We further investigate the correlation in conditional entropy between qubits for different inferred eigenvalues.
-It requires mild anti-correlation for the first two largest eigenvalues and strong anti-correlation for the smallest eigenvalues.
-Most admissible ansatzes lie in a mild anti-correlation regime, explaining VQSD’s limitation in finding the smallest eigenvalues.
-III Main results
+It requires mild correlation for the first two largest eigenvalues and mild anti-correlation for the smallest eigenvalues.
+Most admissible ansatzes lie in a mild correlation regime, explaining VQSD’s limitation in finding the smallest eigenvalues.
+III The reinforcement learning enhanced variational quantum state diagonalisation
+The VQSD algorithm LaRose et al.
+(2019) for a quantum state $\rho$ comprises three subroutines: (1) In the training subroutine, the parameters $\vec{\theta}$ of a quantum gate sequence $U(\vec{\theta})$ are optimised for a given state $\rho$.
+Ideally, after optimisation, the sequence $U(\vec{\theta}_{\text{opt}})$ satisfies
+$\rho^{\prime}=U(\vec{\theta}_{\text{opt}})\rho U(\vec{\theta}_{\text{opt}})^{\dagger}=\rho_{\text{diag}},$
+where $\rho_{\text{diag}}$ is $\rho$ diagonalised in its eigenbasis and $\vec{\theta}_{\text{opt}}$ are the optimal angles.
+Classical gradient-based methods such as SPSA and Gradient Descent, or gradient-free optimisation Powell (1994, 2006) can be utilised in the training process.
+(2) In the eigenvalue readout subroutine, using the optimised unitary $U(\vec{\theta}_{\text{opt}})$ and one copy of state $\rho$, one can extract all the eigenvalues for low-rank states or the largest eigenvalues for full-rank states.
+This is achieved by measuring $\rho^{\prime}$ in the computational basis, $\mathbf{b}=b_{1}b_{2}\ldots b_{n}$, as follows
+$\lambda^{\prime}=\bra{\mathbf{b}}\rho^{\prime}\ket{\mathbf{b}},$
+where $\lambda^{\prime}$ are inferred eigenvalues.
+(3) In the final step, we prepare the eigenvectors associated with the largest eigenvalues.
+If $\mathbf{b}^{\prime}$ is a bit string associated with $\lambda^{\prime}$, then the inferred eigenvectors $|v_{\mathbf{b}^{\prime}}^{\prime}\rangle$ are obtained as follows
+$|{v}^{\prime}_{{\mathbf{b}}^{\prime}}\rangle=U(\theta_{\textrm{opt}})^{\dagger}\ket{{\mathbf{b}}^{\prime}}=U(\theta_{\textrm{opt}})^{\dagger}\left(X^{b_{1}}\otimes\ldots\otimes X^{b_{n}}\right)\ket{\mathbf{0}}.$
+In the reinforcement learning enhanced variational quantum state diagonalisation, an RL-agent is utilised to find the optimal configuration of $U(\vec{\theta})$.
+The agent contains a double deep-Q network (DDQN) Van Hasselt et al.
+(2016) with $\epsilon$-greedy policy.
+The DDQN optimises the stability of Q-learning by using two networks to decouple the action selection $a^{\prime}$ in the subsequent state $s^{\prime}$ from its current reward $r$ evaluation, leading to less overestimated Q-values and more stable and reliable learning outcomes.
+The key update equation is
+$Q(s,a;\theta)\leftarrow Q(s,a;\theta)\\ +\alpha[r+\gamma Q(s^{\prime},\arg\max_{a^{\prime}}Q(s^{\prime},a^{\prime};\theta);\theta^{\prime})-Q(s,a;\theta)]$ (1)
+where, $Q(s,a;\theta)$ is the predicted Q-value for state $s$ and action $a$, parameterised by weights $\theta$ of the online network.
+$\alpha$ is the learning rate and $\gamma$ is the discount factor.
+The settings of the parameters are similar to the ref.
+Kundu et al.
+(2024).
+The optimal policy is obtained by optimising the weights of the DDQN using the ADAM optimiser.
+In the RL setup, the RL-state encodes the ansatz and is defined by a tensor-based encoding presented introduced in Patel et al.
+(2024).
+After each environment and agent interaction, the reward function
+$R=\left\{\begin{array}[]{ll}+\mathcal{R}&\text{for }C_{t}(\vec{\theta})<\zeta+10^{-5}\\ -\textrm{log}\left(C_{t}(\vec{\theta})-\zeta\right)&\text{for }C_{t}(\vec{\theta})>\zeta\end{array}\right.,$ (2)
+is calculated.
+The goal of the RL-agent is to reach the minimum error for a predefined threshold $\zeta$ (which is a hyperparameter of the model) by optimising the parameters of the $U(\vec{\theta})$ using the COBYLA optimiser.
+The cost function
+$C_{t}(\vec{\theta})=\text{Tr}(\rho^{2})-\text{Tr}(\rho_{t}(\vec{\theta})^{2}).$ (3)
+at each step $t$ is calculated for the ansatz which outputs a state $\rho_{t}(\vec{\theta})$.
+IV Analysis of the Haar random input quantum states
+In this section, we analyse the 2-qubit Haar random quantum states.
+The Haar random states can be efficiently simulated using polynomial-time algorithms, as demonstrated in prior research Alagic et al.
+(2020).
+This holds even in scenarios where adversaries have unbounded capabilities with black-box oracle access Chen et al.
+(2024).
+Here, we sample $100000$ Haar random quantum states and observe how the eigenvalues and the qubits’ conditional entropy change with the state’s entanglement.
+Figure 1: The difference in conditional entropy converges towards the zero as the entanglement of the sampled input state increases.
+Meanwhile, as the input state gets closer to the maximally entangled state (with concurrence $1$), the magnitude of the largest eigenvalue increases while the other eigenvalue decreases.
+In Fig.
+1, see that the difference in conditional entropy in qubits of the input state varies evenly on either side of $S_{AB}^{\rho_{AB}}=0$ where $A$ and $B$ are subsystems containing one qubit each and $\rho_{AB}$ is the sampled Haar random state.
+Meanwhile, the area of the distribution of $S_{AB}^{\rho_{AB}}$ on either side of zero entropy converges towards zero as the input state entanglement increases.
+Furthermore, the magnitude of the input quantum state’s largest eigenvalue increases with the increase in entanglement, and all the other eigenvalues converge towards 00.
+It also shows that even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low, justifying the low density of points after $0.6$ in Fig.
+2.
+V Main results
 In this section, we investigate the upper and lower bounds of concurrence of the ansatzes produced by the RL agent in diagonalising $2$-qubit random quantum states sampled from IBM qiskit’s random_density_matrix module.
-For the simulation, we consider multiple such quantum states (specified in the caption of figures), and for each state, we use $10000$ episodes of RL-VQSD to diagonalise the state.
-The configuration of the RL-VQSD is provided in detail in Appendix B.
-III.1 Entanglement
-III.1.1 DDQN weight independent entanglement bound
-To benchmark the performance of the RL-agent for a specific problem, it is ideal to take multiple initialisations of the DDQN (see Appendix B for a brief description).
+For the simulation, we consider multiple such quantum states (specified in the caption of figures), and for each state, we use $10000$ episodes of RL-VQSD algorithm to diagonalise the state.
+The algorithm consists of DDQN with $\epsilon$-greedy policy where the $\epsilon$ decays in each step by a factor of $0.99995$ from its initial value $1$, down to a minimal value $0.05$.
+Furthermore, the network is trained by ADAM optimizer Kingma and Ba (2014) with a learning rate of 0.0001.
+The neural network consists of 5 hidden layers with 1000 neurons and a buffer size of 15000.
+The configuration of the RL-VQSD is provided in detail in Section III.
+V.1 Entanglement
+V.1.1 DDQN weight independent entanglement bound
+To benchmark the performance of the RL-agent for a specific problem, it is ideal to take multiple initialisations of the DDQN (see Section III for a brief description).
 state max. concurrence min. concurrence
@@
 (min.) concurrence is attributed to the ansatz that generates a state with the highest (lowest) concurrence following the evolution of the random quantum state.
-To efficiently investigate the upper and lower bounds of concurrence Hill and Wootters (1997); Wootters (1998) of the admissible ansatz, we calculate the concurrence of admissible ansatz, which is defined by the entanglement induced by the admissible ansatz when initialised with |$\ket{00}$.
+To efficiently investigate the upper and lower bounds of concurrence Hill and Wootters (1997); Wootters (1998) of the admissible ansatz, we calculate the concurrence of the admissible ansatz, which is defined by the entanglement induced by the admissible ansatz when initialised with |$\ket{00}$.
 We first sample different random quantum states, then diagonalise them using RL-VQSD Kundu et al.
@@
 It should be noted that finding the concurrence of the evolved state indicates how much the RL-ansatz induces entanglement to the input sampled state.
-See Appendix C for further analysis of the sampled random quantum states, where we show that in qiskit, random_density_matrix even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low.
+See Section IV for further analysis of the sampled random quantum states, where we show that in qiskit, random_density_matrix even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low.
 The results are presented in Table 1.
@@
 In the following subsection, we investigate how the upper and lower bounds on the entanglement of the RL-ansatz depend on the entanglement of the input quantum state.
-III.1.2 Anti-correlation between entanglement bounds
+V.1.2 Anti-correlation between entanglement bounds
 Corresponding to a predefined threshold, the agent proposes different architectures of quantum circuits that diagonalise the quantum state.
@@
 There is a phase transition from weak anti-correlation to weak correlation.
-In Figure 1, we illustrate the variation of the entanglement bounds with the entanglement of the input quantum state.
+In Fig.
+2, we illustrate the variation of the entanglement bounds with the entanglement of the input quantum state.
 To investigate the dependency, we use the Pearson correlation coefficient (PCC) Lee Rodgers and Nicewander (1988) as a quantifier of correlation.
-See Appendix A for a brief description.
-We see that the upper and lower bounds of entanglement are initially strongly anti-correlation when the input state entanglement is below $0.322$ and then strongly correlate afterwards, indicating a phase change in the upper and lower bounds of entanglement.
+See Section A for a brief description.
+We see that the upper and lower bounds of entanglement are initially strongly anti-correlation when the input state entanglement is below $0.322$ and then strongly correlate afterwards, indicating a phase transition in the upper and lower bounds of entanglement.
 Figure 3: We present the phase transition of correlation between the upper and lower bounds of concurrence for the admissible RL-VQSD ansatz.
@@
 The phase transition point is observed at $k=0.322$ where the concurrence bounds change from mild anti-correlation to mild correlation for the interval $\mathbbm{u}_{2}$ while remaining strongly anti-correlated in the interval $\mathbbm{u}_{1}$.
+Intuitively, when the entanglement of the starting state is low i.e.
+$k\leq 0.322$, it becomes easier for the dephasing subroutine of the VQSD algorithm to eliminate the off-diagonal terms.
+Hence the agent proposes an admissible ansatz to either convert the state into a separable state (eliminating the need for a dephasing gate at the end) or increase its entanglement to the maximum level, thereby making full use of the dephasing gate’s potential.
+This dynamic explains the observed anti-correlation between the entanglement bounds.
+As noted previously, when the entanglement of the starting state is high (in the region $k\geq 0.322$), it becomes easier for the dephasing subroutine to eliminate the off-diagonal terms.
+Thus for states with high entanglement, the agent proposes an admissible ansatz to maximize the entanglement of the starting state.
+Hence we observe the positive relation between the entanglement bounds.
 Next, we obtain the variation of PCC between the upper and lower bound on the concurrence of the initial state.
@@
 We leave proving $\eta^{k}_{w_{ik}}$ is a valid measure of correlation for arbitrary values of $w_{ik}\in(0,1)$ as an open problem.
-We present in Figure 2, the variation of $\eta_{0}^{k}$ (in blue) and $\eta_{1}^{k}$ (in red) for different values of $k$.
+We present in Fig.
+3, the variation of $\eta_{0}^{k}$ (in blue) and $\eta_{1}^{k}$ (in red) for different values of $k$.
 We observe that $\eta_{1}^{k}\approx-1~{}\forall k\in(i,j)$.
@@
 While for values of $k>k_{\ast}$, there is (anti-)correlation between the upper and lower bounds of concurrence of the ansatz for both the intervals $(\mathbbm{u}_{2})~{}\mathbbm{u}_{1}$.
-III.1.3 Enhancing the performance of agent based on initial state entanglement
+The investigation in Fig.
+2 and Fig.
+3 suggests that the structure of the quantum state in terms of its entanglement properties can significantly influence the effectiveness of certain ansatz strategies.
+Specifically, when the entanglement is low i.e.
+below $0.322$, the variability in the entanglement bounds can be exploited.
+For wide entanglement bounds, the flexibility of the ansatz allows for the generation of states that are either highly entangled or nearly separable, starting from $|0\rangle^{\otimes 2}$.
+V.1.3 Enhancing the performance of agent based on initial state entanglement
 Here, we provide a detailed investigation of the admissible ansatzes produced by the RL-agent in terms of ansatz depth, the total number of 2-qubit gates and the total number of gates in the upper and lower bounds of concurrence.
-Figure 3: The RL-agent can generate an ansatz with strong entanglement with fewer 2-qubit gates and circuit depth compared to the ansatzes with weak entanglement.
-As the main aim of RL-VQSD is to diagonalise unitary with a small number of gates and depth, our analysis shows that if we particularly focus on admissible ansatzes near the upper bound of concurrence, we can not only reduce the computational cost of RL-VQSD severely but find optimal admissible ansatzes.
-The numbers in the bars correspond to the amount of entanglement generated by the RL-ansatz when initialised in the vacuum state.
+Figure 4: The RL-agent can generate an admissible ansatz with strong entanglement with fewer 2-qubit gates (2-qubit gate) and circuit depth (Depth) and total number of gates (Numb.
+gate) compared to the ansatzes with weak entanglement.
+As the main aim of RL-VQSD is to diagonalise unitary with a small number of gates and depth, our analysis shows that if we particularly focus on the admissible ansatzes near the upper bound of concurrence, we can not only reduce the computational cost of RL-VQSD severely but also find the optimal admissible ansatzes.
 Our investigation reveals that the admissible ansatzes corresponding to the upper bound of concurrence has, on average, a shorter depth of $13.23$ and a smaller number of gates of $18.5$ compared to the admissible ansatzes corresponding to the lower bound of concurrence where the depth is $13.9$ and number of gates is $19$ over $25$ random quantum states.
-Moreover, Figure 3 makes it more prominent that the minimum number of gates and the depth of the admissible ansatzes are smaller for the upper bound of concurrence than the lower.
+Moreover, Fig.
+4 makes it more prominent that the minimum number of gates and the depth of the admissible ansatzes are smaller for the upper bound of concurrence than the lower.
 This observation is crucial in reducing the subspace of all possible admissible ansatzes.
@@
 In RL-VQSD Kundu et al.
-(2024), the RL-state that encodes the ansatz initiates from an empty circuit (i.e., in circuit 2 $U_{\text{RL-agent}}=\mathds{1}$).
+(2024), the RL-state that encodes the ansatz initiates from an empty circuit (i.e., in circuit 5 the Entanglement Enhancing block is absent and $U_{\text{RL-agent}}=\mathds{1}$).
 We showed that the optimal admissible ansatz, on average, lies in the upper bound of concurrence, motivating us to initialise the quantum state (that is diagonalised using RL-VQSD) closest to the upper bound of entanglement.
@@
 […]
-where the first, Entanglement Enhancing (EE) block, helps the diagonalising state to achieve the highest entanglement before it is fed to the second part, $U_{\text{RL-agent}}$, which is decided by the RL-agent.
-For example, for seed $27$ (see Appendix C for details), we get a random quantum state with concurrence $0.198$, and using the EE, we can increase the input state concurrence to $0.215$ at $\theta=0.5$ as shown in Fig.
-4.
+where the first, Entanglement Enhancing (EE) block, helps the diagonalising state to achieve the highest entanglement before it is fed to the second part, $U_{\text{RL-agent}}$, which the RL-agent decides.
+For example, for seed $27$ (see Section IV for details), we get a random quantum state with concurrence $0.198$ (for seed $27$, and see Appendix A.4 for details on how to generate Haar random state), and using the EE, we can increase the input state concurrence to $0.215$ at $\theta=0.5$ as shown in Fig.
+5.
 Figure 5: The concurrence of input state with $\theta$ of CRX($\theta$) gate in Entanglement Enhancing increases the concurrence of the input random state from $0.198$ to $0.215$ when $\theta=0.5$.
@@
 2 where we see that the number of successful episodes and the reward acquired (where the reward function is in the form defined in Eq.
-12) by the agent in the RL-VQSD process increase by $2$-fold while the EE block is activated for $\theta=0.5$ compared to default settings at $\theta=0$.
+2) by the agent in the RL-VQSD process increase by $2$-fold while the EE block is activated for $\theta=0.5$ compared to default settings at $\theta=0$.
 Table 2: Turning on the Entanglement Enhancing (EE) mode (denoted as Enhanced ($\theta=0.5$)) enhances the learning of the performance of the RL-agent in the absence of the EE mode (denoted as Default ($\theta=0$)).
-In the first $3500$, the Enhanced ($\theta=0.5$) collects $2.08$ times more reward (tot.
+In the first $3500$ episodes, the Enhanced ($\theta=0.5$) collects $2.08$ times more reward (tot.
 rwd.
@@
 Default ($\theta=0$) $6.08\times 10^{-7}$ $11$ $5$
+It should be noted that due to the sake of the interpretability of the EE block, we chose the block resembling a maximally entangled circuit for a 2-qubit system.
+Where the entangler (i.e.
+the CX) gate is parameterised and can induce no entanglement (for $\theta=0$) and maximal entanglement (for $\theta=0.5$) to $|00\rangle$ state.
+The second Hadamard (H) gate on the first qubit is introduced to cancel out the first H in the absence of the entangler at $\theta=0$.
+We can automate the search for gateset in the EE block utilising an RL-agent and diagonalise the input state simultaneously by decomposing the RLQAS into two subroutines.
+Suppose the total number of steps in an episode is $S$, then in the first subroutine for $s_{\text{EE}}<S$ steps the agent’s task would be to find a gateset (consists of maximum $s_{\text{EE}}$ gates) that maximises the entanglement of the input state by utilising a reward function $R_{\text{EE}}$.
+One can formulate $R_{\text{EE}}$ similar to ref.
+Kuo et al.
+(2021).
+Wherein in the second subroutine, for the remaining steps $S-s_{\text{EE}}$, the agent focuses on finding possible structures of $U_{\text{RL-ansatz}}$ to diagonalise the output state of the EE block using the reward defined in Eq.
+2.
 The aforementioned observations lead us to the conclusion that maximising the entanglement of the input state through the initialisation strategy of Fig.
 5 enhances the learning process of the RL-agent and facilitates the further optimisation of the admissible ansatzes compared to the start-of-art Kundu (2024).
-III.2 Conditional quantum entropy
-III.2.1 Contribution of individual qubits for change in entanglement
+V.2 Conditional quantum entropy
+V.2.1 Contribution of individual qubits for change in entanglement
 We quantify the contribution of each qubit using the conditional quantum entropy $S_{q_{0}|q_{1}}^{\rho^{\prime}_{q_{0}q_{1}}}$ Cerf and Adami (1997, 1999); Wilde (2013).
@@
 Hence, it is essential to know how much information the state of qubits provides about the state of the $N$ qubits.
-For a detailed discussion and notation, see Appendix A.
-Figure 5: Relative change in the contribution of qubits 0 and 1 as a function of change in entanglement.
-We quantify the relative contribution of individual qubits to the RL-VQSD ansatz using $\Delta_{q_{0}}^{q_{1}}$ (see Eq.
+For a detailed discussion and notation, see Section A.
+Figure 6: We analyze the relative change in the contribution of the first and second qubits as a function of changes in entanglement.
+To quantify the relative contribution of individual qubits to the RL-VQSD ansatz, we use $\Delta_{q_{0}}^{q_{1}}$ (see Eq.
 (6)).
 For the change in concurrence due to the ansatz, the distribution of $\Delta_{q_{0}}^{q_{1}}$ is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$.
-The magnitude of the cumulative weight of all data points in the region $\Delta_{q_{0}}^{q_{1}}<0$ is $4.854$ while that for data points in the region $\Delta_{q_{0}}^{q_{1}}>0$ is $5.524$.
-This indicates equal contribution from both the qubits to the RL-VQSD ansatz for the range of change in entanglement.
+The cumulative weight of all data points in the region $\Delta_{q_{0}}^{q_{1}}<0$ is 4.854, while that for data points in the region $\Delta_{q_{0}}^{q_{1}}>0$ is 5.524.
+This indicates an equal contribution from both qubits to the RL-VQSD ansatz across the range of changes in entanglement.
 Let the change in the conditional entropy of the qubit $q_{0}(q_{1})$ after evolving through the VQSD ansatz be given by $\Delta S^{\rho\to{\rho^{\prime}}}_{q_{0}|q_{1}}(\Delta S^{\rho\to{\rho^{\prime}}}_{q_{1}|q_{0}})$.
@@
 Following a similar reasoning $\Delta_{q_{0}}^{q_{1}}<0$ implies $q_{1}$ has greater contribution as compared to $q_{0}$, while $\Delta_{q_{0}}^{q_{1}}=0$ implies both $q_{0}$ and $q_{1}$ contributes equally to the ansatz.
-We present the variation of $\Delta_{q_{0}}^{q_{1}}$ for the change in concurrence due to the ansatz given by $\Delta_{c}\coloneqq C^{\rho_{q_{0}q_{1}}}-C^{\rho^{\prime}_{q_{0}q_{1}}}$ in Figure 5.
+We present the variation of $\Delta_{q_{0}}^{q_{1}}$ for the change in concurrence due to the ansatz given by $\Delta_{c}\coloneqq C^{\rho_{q_{0}q_{1}}}-C^{\rho^{\prime}_{q_{0}q_{1}}}$ in Fig.
+6.
 We observe that the distribution of $\Delta_{q_{0}}^{q_{1}}$ is bounded in the interval $[-0.7,0.7]$ and is uniform with respect to $\Delta_{q_{0}}^{q_{1}}=0$.
@@
 Also, for $\Delta_{c}<0.2$, $\Delta_{q_{0}}^{q_{1}}$ increases monotonically with increase in $\Delta_{c}$.
-III.2.2 Problem with variational quantum state diagonalisation
+V.2.2 Problem with variational quantum state diagonalisation
 In this section, we investigate the contribution of individual qubits to the performance of the RL-VQSD algorithm.
 Through this investigation, we get a deeper insight into why the variational quantum state diagonalization algorithm can not find the smallest eigenvalues.
-We present the main observations in Figure 6, where we plot the four eigenvalues of $21$, $2$-qubit Haar random mixed quantum states with the correlation in conditional entropy among the qubits.
-For an individual quantum state, the RL-agent proposes $\mathbf{s}$ different structures of admissible ansatzes.
+We present the main observations in Fig.
+7, where we plot the four eigenvalues of $35$, $2$-qubit Haar random mixed quantum states with the correlation in conditional entropy among the qubits.
+For an individual quantum state, the RL-agent proposes $\mathbf{s}$ different structures of the admissible ansatzes.
 Then we get $\rho_{i}^{\prime}$ states where $i\in[0,\mathbf{s}]$.
@@
 Then, we calculate the correction in conditional entropy among the qubits using PCC.
-In Figure 6, we observe that a mild anti-correlation between the qubits is required to find the two largest eigenvalues.
+In Fig.
+7, we observe that a mild anti-correlation between the qubits is required to find the two largest eigenvalues.
 As the magnitude of the largest eigenvalues decreases, the mild anti-correlation turns into a strong correlation between the qubits.
-This indicates a strong correlation between the qubits is required to achieve the two largest eigenvalues.
-Meanwhile, we require a strong anti-correlation between the qubits to find the smallest eigenvalues of the same states.
+This indicates that the qubits require mild to strong correlation to achieve the two largest eigenvalues.
+Meanwhile, we require a mild to strong anti-correlation between the qubits to find the smallest eigenvalues of the same states.
 Then, as the magnitude of the smallest eigenvalue increases, the strong anti-correlation turns into mild anti-correlation.
-Figure 6: The conditional entropy correlation between the first and the second qubit decreases as the magnitude of the largest and the second largest eigenvalues increases.
-Meanwhile, the same correlation between the qubits increases with the magnitude of the third and the smallest eigenvalues.
-The RL-ansatz proposes specific structures of admissible ansatzes, and for each such ansatz, it is not possible to have two different kinds of correlations.
-Hence, our investigation suggests that it is not feasible to find the smallest eigenvalues with the same ansatz that provides a good approximation of the largest eigenvalues.
-The correlation of conditional entropy between qubits is calculated using Pearson correlation coefficient (PCC) and is the median over all the configuration of RL-ansatz proposed by the agent.
+Figure 7: On average over $35$ random quantum states the correlation in conditional entropy between the first and second qubits decreases as the magnitude of the largest and second largest eigenvalues increases.
+Meanwhile, the same correlation between the qubits increases with the magnitude of the third largest and the smallest eigenvalues.
+The RL-ansatz proposes specific structures of admissible ansatzes, and for each such ansatz, it is not possible to have both types of correlations simultaneously.
+Hence, our investigation suggests that it is not feasible to find the smallest eigenvalues using the same ansatz that provides a good approximation of the largest eigenvalues.
+The correlation of conditional entropy between qubits is calculated using the Pearson correlation coefficient (PCC) and is the median over all configurations of RL-ansatz proposed by the agent.
 The $\bigstar$ signifies the eigenvalues corresponding to the same state.
 We recall that the correlation is calculated among the qubits based on the conditional entropy of individual qubits.
-As it is not possible to have two different kinds (such as mild anti-correlation and strong anti-correlation at the same time) of correlation between the qubits in the same ansatz structure, we can not find the largest and smallest eigenvalues utilising the same ansatz.
-For example, we mark the eigenvalues of a state with $\bigstar$, where we observe that to find the largest two eigenvalues, the RL-ansatz need to induce a mild anti-correlation between the two qubits; meanwhile, to obtain the smallest eigenvalues, we require strong anti-correlation.
-On the contrary, for the algorithm, as the agent adds a new gate to the RL-ansatz, the correlation in conditional entropy among the qubits increases.
-That is why it is easier for the RL-ansatz to propose admissible ansatz with mild anti-correlation.
-We further certify this observation in Figure 7, where we observe the correlation in conditional entropy between qubits for $21$ random quantum states.
-We see that most of the admissible ansatz provide a mild anti-correlation among the qubits, indicating that the VQSD algorithm primarily focuses on finding the largest eigenvalues rather than the smallest ones.
-Hence, it is easier for the variational diagonalisation algorithm to find the largest eigenvalues more efficiently than the smallest ones.
-Figure 7: The correlation in conditional entropy among qubits is in the range of mild correlation for $21$ random quantum states.
-This observation, along with the observation in Figure 6 certify that RL-agent only proposes a particular configuration of admissible ansatz that can find only the largest eigenvalues.
-The average correlation among all the $21$ states is $-0.4263$ (horizontal black line), which is in the regime of mild anti-correlation.
-IV Discussion
+As it is impossible to have two different kinds of correlation (such as mild correlation and mild anti-correlation simultaneously) between the qubits in the same ansatz structure, we can not find the largest and smallest eigenvalues utilising the same ansatz.
+For example, we mark the eigenvalues of a state with $\bigstar$, where we observe that to find the largest two eigenvalues, the RL-ansatz need to induce a mild correlation between the two qubits; meanwhile, to obtain the smallest eigenvalues, we require mild anti-correlation.
+Figure 8: The average correlation in conditional entropy among qubits is in the range of mild correlation for $35$ random quantum states for the first two largest eigenvalues whereas the conditional entropy among qubits requires a mild anti-correlation for the smallest eigenvalues.
+This observation, along with the observation in Fig.
+7 certify that with a specific configuration of an admissible ansatz, we can not simultaneously find the largest and the smallest eigenvalues.
+The average correlation in conditional entropy among qubits falls within the range of mild correlation for the first two largest eigenvalues across $35$ random quantum states.
+Conversely, the conditional entropy among qubits exhibits a mild anti-correlation for the smallest eigenvalues.
+This finding, together with the observation in Fig.
+8, indicates that with a specific configuration of an admissible ansatz, it is interactable to simultaneously identify the largest and smallest eigenvalues.
+It should be noted that for the $35$ random quantum states, the average correlation in conditional entropy lies in the mild correlation zone, indicating that RL-VQSD can efficiently find the largest few eigenvalues.
+However, it fails to provide a good approximation for the smallest eigenvalues simultaneously.
+VI Discussion
 In this article, we use tools from quantum information theory to analyse the admissible ansatzes proposed by the RL-agent for solving reinforcement learning-based quantum architecture search problems, and we particularly investigate the RL-agent proposed ansatzes in solving recently proposed RL-VQSD problem.
@@
 The upper and lower bounds are initially anti-correlated with respect to the initial state entanglement.
-The anti-correlation turns into a mild and eventually strong correlation as the entanglement of the initial state surpasses beyond a phase change point at concurrence $0.322$.
+The anti-correlation turns into a mild and eventually strong correlation as the entanglement of the initial state surpasses beyond a phase transition point at concurrence $0.322$.
 We also see that the optimal configuration of the admissible ansatz lies in the upper bound of concurrence, which has a smaller circuit depth and requires fewer gates than the lower bound.
@@
 We effectively utilise these observations to greatly reduce the RL-agent and environment interaction, i.e., the computational time needed to solve the RL-VQSD problem.
-Specifically, we propose an entanglement enhancing initial block that adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$ Additionally, the admissible ansatzes not only diagonalise the random quantum states with high accuracy but can be used to generate close to maximum entangled states from the vacuum state.
+Specifically, we propose an entanglement enhancing block that adjusts the input state’s concurrence thereby enhancing the performance of the RL-agent by a factor of $2$.
+Additionally, the admissible ansatzes not only diagonalise the random quantum states with high accuracy but can be used to generate close to maximum entangled states from the vacuum state.
 Furthermore, we quantify the contribution of each qubit in the admissible ansatz using conditional quantum entropy.
 For the admissible RL-VQSD ansatz, we observe an equal contribution from each qubit for the full range of change in the entanglement of the state due to the ansatz.
-Focusing on the task of obtaining the eigenvalues of the starting state, we observe that a strong correlation between the qubits of the admissible ansatz is required to obtain the two largest eigenvalues, while strong anti-correlation between the qubits is needed to find the smallest eigenvalues of the same state.
-Noting that it is impossible to have both strong correlation and anti-correlation between the qubits of the same ansatz, we conclude that we cannot simultaneously obtain both the largest and smallest eigenvalues from the RL-VQSD admissible ansatz.
-From the observation, most of the admissible ansatz has mild anti-correlation among the qubits, which explains why it is easier for the diagonalisation algorithm to find the largest eigenvalues than the smallest ones.
-For future work, it would be interesting to conduct more comprehensive study of the implication of introducing entanglement Horodecki et al.
-(2009) or nonlocality Brunner et al.
+Focusing on the task of obtaining the eigenvalues of the starting state, we observe that a mild correlation between the qubits of the admissible ansatz is required to obtain the two largest eigenvalues, while mild anti-correlation between the qubits is needed to find the smallest eigenvalues of the same state.
+Noting that it is impossible to have both mild correlation and anti-correlation between the qubits of the same ansatz, we cannot simultaneously obtain both the largest and smallest eigenvalues from the RL-VQSD admissible ansatz.
+From the observation, most of the admissible ansatz has mild correlation among the qubits, which explains why it is easier for the VQSD algorithm to find the largest eigenvalues than the smallest ones.
+For future work, one could conduct a comprehensive study of the implication of introducing correlations such as entanglement Horodecki et al.
+(2009), nonlocality Brunner et al.
+(2014); Sadhu and Das (2023) or steering Uola et al.
 (2020); Sadhu and Das (2024)-based reward functions of the RL-agent in QAS.
 Specifically, given that most of the admissible ansatzes have high values of concurrence, it is expected that an entanglement or non-locality-based reward function may further reduce the time required to find the optimal configurations.
-Also, we may expect this to provide an optimal quantum circuit that provides both the lowest and highest eigenvalues of the input quantum state.
+Also, we may expect this to provide an optimal quantum circuit that provides both the eigenvalues of a random quantum state with higher accuracy within a smaller time.
+In this work, we present an intuitive explanation of the phase transition behavior of the bounds on the entanglement of the admissible ansatz.
+However, we point out that the phase transition point is problem-specific.
+It would be interesting to see if the phase transition behavior still holds for other classes of algorithms, a problem we leave for future research.
+Furthermore, using a similar approach, presented in our paper one can rigorously investigate the RL-ansatz in the task of minimizing the entanglement entropy of a quantum many-body system to identify quantum phase transition An et al.
+(2021) and finding the ground state of chemical hamiltonian Patel et al.
+(2024); Ostaszewski et al.
+(2021).
 Acknowledgements.
@@
 Observation 1.
-For the state $\rho^{\prime}_{AB}\in\mathscr{D}(\mathscr{H}_{A}\otimes\mathscr{H}_{B})$, consider a purification |$\ket{\psi}_{ABE}\in\mathscr{H}_{A}\otimes\mathscr{H}_{B}\mathscr{H}_{E}$.
+For the state $\rho^{\prime}_{AB}\in\mathscr{D}(\mathscr{H}_{A}\otimes\mathscr{H}_{B})$, consider a purification |$\ket{\psi}_{ABE}\in\mathscr{H}_{A}\otimes\mathscr{H}_{B}\otimes\mathscr{H}_{E}$.
 It then follows that:
@@
 The conditional quantum information thus measures the difference in the entropy of the states $\rho_{E}\in\mathscr{D}(\mathscr{H}_{E})$ and $\rho_{B}\in\mathscr{D}(\mathscr{H}_{B})$.
-Theorem 1.
-For the state $\rho^{\prime}_{AB}\in\mathscr{D}(\mathscr{H}_{A}\otimes\mathscr{H}_{B})$, the following inequality applies to the marginal entropy $S_{A}^{\rho^{\prime}_{AB}}$ and the conditional quantum entropy $S_{A|B}^{\rho^{\prime}_{AB}}$:
-$S_{A}^{\rho^{\prime}_{AB}}\geq S_{A|B}^{\rho^{\prime}_{AB}}.$ (8)
-The above theorem states that conditioning cannot increase entropy.
-The value of conditional entropy can be positive or negative depending on the quantum state for which it is evaluated.
-However, it cannot be arbitrarily large or arbitrarily small.
-The following theorem places an upper bound on the absolute value of the conditional quantum entropy.
-Theorem 2.
-The conditional entropy is bounded as follows:
-$|S_{A|B}^{\rho^{\prime}_{AB}}|\leq\log\dim(\mathscr{H}_{A})$ (9)
-The bound is saturated for $\rho^{\prime}_{AB}=\mathbbm{1}_{A}\otimes\sigma_{B}$, where $\mathbbm{1}_{A}$ is the maximally mixed state and $\sigma_{B}\in\mathscr{D}(\mathscr{H}_{B})$, and also for the family of maximally entangled states.
 In the following subsection, we discuss a measure to quantify the amount of correlation between two random variables.
@@
 A correlation value of +1(1)11+1(-1)+ 1 ( - 1 ) implies perfect correlation (anti-correlation) between the two random variables $r_{1}$ and $r_{2}$, while zero implies no linear dependence between the variables.
-Appendix B The reinforcement learning enhanced variational quantum state diagonalisation
-The VQSD algorithm LaRose et al.
-(2019) for a quantum state $\rho$ comprises three subroutines: (1) In the training subroutine, the parameters $\vec{\theta}$ of a quantum gate sequence $U(\vec{\theta})$ are optimised for a given state $\rho$.
-Ideally, after optimisation, the sequence $U(\vec{\theta}_{\text{opt}})$ satisfies
-$\rho^{\prime}=U(\vec{\theta}_{\text{opt}})\rho U(\vec{\theta}_{\text{opt}})^{\dagger}=\rho_{\text{diag}},$
-where $\rho_{\text{diag}}$ is $\rho$ diagonalised in its eigenbasis and $\vec{\theta}_{\text{opt}}$ are the optimal angles.
-Classical gradient-based methods such as SPSA and Gradient Descent, or gradient-free optimisation Powell (1994, 2006), can be utilised in the training process.
-(2) In the eigenvalue readout subroutine, using the optimised unitary $U(\vec{\theta}_{\text{opt}})$ and one copy of state $\rho$, one can extract all the eigenvalues for low-rank states or the largest eigenvalues for full-rank states.
-This is achieved by measuring $\rho^{\prime}$ in the computational basis, $\mathbf{b}=b_{1}b_{2}\ldots b_{n}$, as follows
-$\lambda^{\prime}=\bra{\mathbf{b}}\rho^{\prime}\ket{\mathbf{b}},$
-where $\lambda^{\prime}$ are inferred eigenvalues.
-(3) In the final step, we prepare the eigenvectors associated with the largest eigenvalues.
-If $\mathbf{b}^{\prime}$ is a bit string associated with $\lambda^{\prime}$, then the inferred eigenvectors $|v_{\mathbf{b}^{\prime}}^{\prime}\rangle$ are obtained as follows
-$|{v}^{\prime}_{{\mathbf{b}}^{\prime}}\rangle=U(\theta_{\textrm{opt}})^{\dagger}\ket{{\mathbf{b}}^{\prime}}=U(\theta_{\textrm{opt}})^{\dagger}\left(X^{b_{1}}\otimes\ldots\otimes X^{b_{n}}\right)\ket{\mathbf{0}}.$
-In the reinforcement learning enhanced variational quantum state diagonalisation, an RL-agent is utilised to find the optimal configuration of $U(\vec{\theta})$.
-The agent contains a double deep-Q network (DDQN) Van Hasselt et al.
-(2016) with $\epsilon$-greedy policy.
-The DDQN optimises the stability of Q-learning by using two networks to decouple the action selection $a^{\prime}$ in the subsequent state $s^{\prime}$ from its current reward $r$ evaluation, leading to less overestimated Q-values and more stable and reliable learning outcomes.
-The key update equation is
-$Q(s,a;\theta)\leftarrow Q(s,a;\theta)\\ +\alpha[r+\gamma Q(s^{\prime},\arg\max_{a^{\prime}}Q(s^{\prime},a^{\prime};\theta);\theta^{\prime})-Q(s,a;\theta)]$ (11)
-where, $Q(s,a;\theta)$ is the predicted Q-value for state $s$ and action $a$, parameterised by weights $\theta$ of the online network.
-$\alpha$ is the learning rate and $\gamma$ is the discount factor.
-The settings of the parameters are similar to the ref.
-Kundu et al.
-(2024).
-The optimal policy is obtained by optimising the weights of the DDQN using the ADAM optimiser.
-In the RL setup, the RL-state encodes the ansatz and is defined by a tensor-based encoding presented introduced in Patel et al.
-(2024).
-After each environment and agent interaction, the reward function
-$R=\left\{\begin{array}[]{ll}+\mathcal{R}&\text{for }C_{t}(\vec{\theta})<\zeta+10^{-5}\\ -\textrm{log}\left(C_{t}(\vec{\theta})-\zeta\right)&\text{for }C_{t}(\vec{\theta})>\zeta\end{array}\right.,$ (12)
-is calculated.
-The goal of the RL-agent is to reach the minimum error for a predefined threshold $\zeta$ (which is a hyperparameter of the model) by optimising the parameters of the $U(\vec{\theta})$ using the COBYLA optimiser.
-The cost function
-$C_{t}(\vec{\theta})=\text{Tr}(\rho^{2})-\text{Tr}(\rho_{t}(\vec{\theta})^{2}).$ (13)
-at each step $t$ is calculated for the ansatz which outputs a state $\rho_{t}(\vec{\theta})$.
-Appendix C Analysis of the Haar random input quantum states
-In this section, we analyse the 2-qubit quantum states sampled from random_density_matrix of IBM qiskit’s quantum_info module.
-Here, we sample $100000$ Haar random quantum states and observe how the eigenvalues and the qubits’ conditional entropy change with the state’s entanglement.
-Figure 8: The difference in conditional entropy converges towards the zero as the entanglement of the sampled input state increases.
-Meanwhile, as the input state gets closer to the maximally entangled state (with concurrence $1$), the magnitude of the largest eigenvalue increases while the other eigenvalue decreases.
-In Fig.
-8, see that the difference in conditional entropy in qubits of the input state varies evenly on either side of $S_{AB}^{\rho_{AB}}=0$ where $A$ and $B$ are subsystems containing one qubit each and $\rho_{AB}$ is the sampled Haar random state.
-Meanwhile, the area of the distribution of $S_{AB}^{\rho_{AB}}$ on the ether side of zero entropy converges towards zero as the input state entanglement increases.
-Furthermore, the magnitude of the input quantum state’s largest eigenvalue increases with the increase in entanglement, and all the other eigenvalues converge towards 00.
-It also shows that even for $100000$ seeds, the probability of sampling states with concurrence greater than $0.6$ is very low, justifying the low density of points after $0.6$ in Fig.
-1.
-Sampling a random quantum state
-If $n$ is the number of qubits and $s$ is the random seed to sample, then we can sample a state from random_density_matrix of IBM qiskit’s quantum_info module (in python) as follows
+A.4 Sampling a random quantum state
+If $n$ is the number of qubits and $s$ is the random seed to sample, then we can sample a state from random_density_matrix of IBM qiskit’s quantum_info module as follows
 […]
-The state used in III.1, is sampled using $n=2$ and $s=27$
+The state used in V.1, is sampled using $n=2$ and $s=27$.